### v1

# The tensor product of $p$-adic Hilbert spaces

Abstract
In the framework of quantum mechanics over a quadratic extension of the ultrametric field of $p$-adic numbers, we introduce a notion of tensor product of $p$-adic Hilbert spaces. To this end, following a standard approach, we first consider the algebraic tensor product of $p$-adic Hilbert spaces. We next define a suitable norm on this linear space. It turns out that, in the $p$-adic framework, this norm is the analogue of the projective norm associated with the tensor product of real or complex normed spaces. Eventually, by metrically completing the resulting $p$-adic normed space, and equipping it with a suitable inner product, we obtain the tensor product of $p$-adic Hilbert spaces. That this is indeed the correct $p$-adic counterpart of the tensor product of complex Hilbert spaces is also certified by establishing a natural isomorphism between this $p$-adic Hilbert space and the corresponding Hilbert-Schmidt class. Since the notion of subspace of a $p$-adic Hilbert space is highly nontrivial, we finally study the tensor product of subspaces, stressing both the analogies and the significant differences with respect to the standard complex case. These findings should provide us with the mathematical foundations necessary to explore quantum entanglement in the $p$-adic setting, with potential applications in the emerging field of $p$-adic quantum information theory.
1 Introduction
During the final part of the 20th century, $p$-adic mathematical physics emerged as an effort to elaborate a non-Archimedean approach to space-time at very small scales. In fact, it is a general belief that standard quantum mechanics over the field of complex numbers is no more applicable below a scale of the order of Planck’s length ($l_{\text{P}}=\sqrt{\hbar G/c^{3}}\approx 10^{-35}\text{m}$), which is conjectured to be the smallest measurable length [1]. This belief led some theoretical physicists to consider the formulation of a non-Archimedean quantum physics as a viable description of quantum gravity and string theory [2, 3, 4, 5]. The adoption of the field $\mathbb{Q}_{p}$ of $p$-adic numbers, where $p$ is a prime number, is strongly favored because it is, up to isomorphisms, the only complete non-Archimedean field obtainable by metrically extending rational numbers [6, 7].
According to the models proposed so far, the field $\mathbb{Q}_{p}$ can be invoked at two different levels when formulating a $p$-adic quantum mechanical theory. On the one hand, one may simply resort to wave functions of the form $\psi\colon\mathbb{Q}_{p}\to\mathbb{C}$ (or, more generally, complex wave functions on a $p$-adic manifold), in such a way to incorporate the non-Archimedean behavior of the theory directly into the space-time coordinates [8, 9, 10, 11, 12, 13, 14, 15]. On the other hand, $\mathbb{Q}_{p}$ — or rather, for technical reasons related, e.g., to the role played by a suitable conjugation map in the theory, a quadratic extension of this field — may be adopted as the base field of the carrier vector space of physical states [16, 17, 18, 19, 20, 21]. This is both a more subtle and more significant modification of the standard formulation of quantum mechanics that entails, among other aspects, the development of new mathematical tools and a suitable interpretation of the physically observable quantities [22, 23, 24].
In a preceding paper [22], a new approach to quantum mechanics over the non-Archimedean field of $p$-adic numbers has been proposed that mimics, to some extent, the algebraic formulation of standard quantum mechanics [25]. The main novelty of this approach lies in a suitable definition of a $p$-adic Hilbert space and its remarkable consequences. A central aspect of our definition of a $p$-adic Hilbert space is that the existence of orthonormal bases is part of the fundamental axioms rather than an effect of these axioms. This peculiar fact is due to the very special interplay between the non-Archimedean norm and the scalar product [22], that has no parallel in the standard complex case.
As previously mentioned, in our formulation of $p$-adic quantum mechanics, physical states are defined as linear functionals on a *-algebra of observables over a quadratic extension $\mathbb{Q}_{p,\mu}$ of the field of $p$-adic numbers, where $\mu$ is a parameter labeling the various possible quadratic extensions of $\mathbb{Q}_{p}$ (for generalities regarding the quadratic extensions of $p$-adic numbers, see [6, 7]). At this stage, as pointed out in [22], an important step forward in developing a complete theory of $p$-adic quantum mechanics is to define the tensor product of two $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$, to provide a mathematical description of a composite physical system. In our present contribution, we address this issue precisely.
Let us briefly outline our strategy. First, following the usual approach, we define the algebraic tensor product [26] $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ of two $p$-adic Hilbert spaces, which amounts to considering them as mere linear spaces over the field $\mathbb{Q}_{p,\mu}$. Next, a non-Archimedean norm is introduced, suitable for achieving a $p$-adic normed space $\mathcal{H}\otimes_{\pi}\mathcal{K}$ out of the initial algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}$. This norm may be regarded as the ultrametric analogue of the standard projective norm [26] associated with the algebraic tensor product of two real or complex Banach spaces — which explains the meaning of the symbol $\mathcal{H}\otimes_{\pi}\mathcal{K}$ we adopt here — and, accordingly, the associated Banach space completion will be denoted by $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. Eventually, by endowing the $p$-adic Banach space $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$ with a suitable inner product, a $p$-adic Hilbert space is obtained; i.e., the ($p$-adic Hilbert space) tensor product $\mathcal{H}\otimes\mathcal{K}$ of the original Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$. In particular, an orthonormal basis is explicitly constructed in such a way as to satisfy a fundamental axiom of a $p$-adic Hilbert space (i.e., the existence of such a basis). We also show that the tensor product $p$-adic Hilbert space $\mathcal{H}\otimes\mathcal{K}$ is isomorphic to the space of trace class operators — which in turn, endowed with a suitable scalar product, may be identified with the space of Hilbert-Schmidt operators (the so-called $p$-adic trace/Hilbert-Schmidt class) — from $\mathcal{H}$ into $\mathcal{K}$. Obtaining this result may be regarded as a sort of litmus test that our construction (which, differently form the standard complex setting, involves the projective tensor product) provides indeed the correct $p$-adic analogue of the usual tensor product of two complex Hilbert spaces.
Finally, we study how the tensor product $\mathcal{H}\otimes\mathcal{K}$ behaves w.r.t. the subspaces of the original $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$. In this regard, among many other technical issues, a fundamental point that one should keep in mind is the following. The notion of subspace itself is not as simple and straightforward in the $p$-adic setting as it is in the complex case [23]. Accordingly, the analogies — as well as several nontrivial differences — with respect to the standard complex setting, are analyzed throughout the article.
The paper is structured as follows. In Section 2, we recall the main definitions and properties of $p$-adic Hilbert spaces, as well as some basic facts about linear operators between $p$-adic Hilbert spaces that are relevant for our purposes. Section 3 is the core part of the paper, since it is devoted to the construction of the tensor product of two $p$-adic Hilbert spaces, along the lines traced above. Next, in Section 4, we prove the aforementioned isomorphism between the tensor product $\mathcal{H}\otimes\mathcal{K}$ and the $p$-adic trace/Hilbert-Schmidt class of operators from $\mathcal{H}$ into $\mathcal{K}$. The implementation of this isomorphism involves a notion of anti-unitary operator suitable for the $p$-adic framework. In Section 5, we analyze the properties of the tensor product of two $p$-adic Hilbert spaces w.r.t. the subspaces of the two factor spaces. Finally, in Section 6, conclusions are drawn, and we take a look at some research prospects.
2 Preliminaries
In this section, we collect some basic results and tools that will be relevant for our later purposes. To begin with, we discuss the $p$-adic normed and Hilbert spaces over a quadratic extension $\mathbb{Q}_{p,\mu}$ of the field $\mathbb{Q}_{p}$ of $p$-adic numbers (for a comprehensive discussion of $\mathbb{Q}_{p}$ and its quadratic extensions, the reader may refer, e.g., to [6, 7]). We next focus on the classes of bounded, adjointable, and trace class/Hilbert-Schmidt operators between $p$-adic Hilbert spaces.
Even if the reader may find a more detailed account in the preceding papers [22, 23, 24] — where, for the sake of simplicity, linear operators defined on a single Hilbert space are considered — it is worth discussing here how the main results can be extended to operators acting, in general, between two different $p$-adic Hilbert spaces — say, $\mathcal{H}$ and $\mathcal{K}$ — as this generalization will provide us with an essential tool for studying, in Section 4, the relation between the $p$-adic trace/Hilbert-Schmidt class operators from $\mathcal{H}$ to $\mathcal{K}$ and the tensor product $\mathcal{H}\otimes\mathcal{K}$ of the two Hilbert spaces.
2.1 $p$-adic Hilbert spaces
We recall that a $p$-adic normed space is a pair $(X,\|\cdot\|)$, where $X$ is a vector space over $\mathbb{Q}_{p,\mu}$, and $\|\cdot\|$ an ultrametric (or non-Archimedean) norm defined on $X$, namely, a map $\|\cdot\|\colon X\rightarrow\mathbb{R}^{+}$ that is positive definite, absolute-homogeneous and satisfies the so-called strong triangle inequality, i.e.,
$\|x+y\|\leq\max(\|x\|\,,\|y\|),\quad\forall x,y\in X.$ (1)
We say that $\|\cdot\|$ is an ultrametric seminorm if it is absolute-homogeneous and satisfies (1). If $X$ is complete w.r.t. the (ultra)-metric induced by $\|\cdot\|$, the pair $(X,\|\cdot\|)$ is called a $p$-adic Banach space [7, 22, 27, 23, 29, 30, 31]. In the following, we will consider separable $p$-adic Banach spaces.
Definition 2.1.
Let $(X,\|\cdot\|)$ be a $p$-adic normed space. We say that two vectors $x,y$ in $X$ are (mutually) norm-orthogonal — and we write $x\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}y$ — if the condition
$\|\alpha x+\beta y\|=\max(\|\alpha x\|\,,\|\beta y\|)$ (2)
holds for every $\alpha,\beta\in\mathbb{Q}_{p,\mu}$. We say that $\{x,y\}\subset X\setminus\{0\}$ is a norm-orthogonal system if $x$ and $y$ are mutually norm-orthogonal; we say that $\{x,y\}$ is a normal system if $x$ and $y$ are mutually norm-orthogonal, and, in addition, they are normalized, i.e., $\|x\|=\|y\|=1$.
Remark 2.2.
The previous notion of orthogonality of vectors is equivalent to the so-called Birkhoff-James orthogonality [32, 33]; i.e., $x\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}y$ iff $\|x\|\leq\|x+\alpha y\|$ for all $\alpha\in\mathbb{Q}_{p,\mu}$ (clearly, in a real or complex Hilbert space this notion is equivalent to the usual one). The geometric intuition behind the latter definition is the following: the vector $x$ is orthogonal to the vector $y$ iff every triangle with one side equal to $x$ and another side parallel to $y$ has its third side that is not shorter than $x$. Note that we have: $0\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}x$, for all $\in X; xx$$\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}x\;\implies\;x=0$; if $x\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}y$, then $\alpha x\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}\beta y$ for all $\alpha,\beta\in\mathbb{Q}_{p,\mu}$.
More generally, given a $p$-adic normed space $(X,\|\cdot\|)$, we say that a (nonempty) set $\mathfrak{A}\subset X\setminus\{0\}$ is a norm-orthogonal system if either $\mathfrak{A}$ is a singleton or, for any $x,y\in\mathfrak{A}$, with $x\neq y$, $\{x,y\}$ is a norm-orthogonal system; we say that $\mathfrak{A}$ is a normal system if either $\mathfrak{A}=\{x\}$, with $\|x\|=1$, or, for any $x,y\in\mathfrak{A}$, with $x\neq y$, $\{x,y\}$ is a normal system. Otherwise stated, a (nonempty) set $\mathfrak{A}\subset X\setminus\{0\}$ is a normal system if it consists of normalized vectors that are pairwise norm-orthogonal.
Definition 2.3.
A countable subset $\mathfrak{B}$ of a (separable) $p$-adic Banach space $(X,\|\cdot\|)$ is called a norm-orthogonal basis (alternatively, a normal basis) if $\mathfrak{B}$ is a norm-orthogonal (respectively, a normal) system and, moreover, if for each $x\in X$ there exists a map $\alpha_{x}\colon\mathfrak{B}\rightarrow\mathbb{Q}_{p,\mu}$ such that $x$ can be expressed — in a unique way, up to reordering of the summands — as the sum of an unconditionally norm-convergent series of the form
$x=\sum_{b\in\mathfrak{B}}\alpha_{x}(b)\hskip 1.70717ptb.$ (3)
Remark 2.4.
Note that a norm-orthogonal basis in a $p$-adic Banach space is an unconditional Schauder basis [22].
One can prove that a (separable) $p$-adic Banach space $(X,\|\cdot\|)$ always admits a norm-orthogonal basis, while a normal basis exists in $X$ iff the additional condition $\|X\|=\left|\mathbb{Q}_{p,\mu}\right|$ is satisfied [7, 22, 29] — where $\|X\|$ and $\left|\mathbb{Q}_{p,\mu}\right|$ are subsets of $\mathbb{R}^{+}$ defined by
$\|X\|\coloneqq\{\|x\|\mid x\in X\},\quad\left|\mathbb{Q}_{p,\mu}\right|\coloneqq\{\left|\alpha\right|\mid\alpha\in\mathbb{Q}_{p,\mu}\}.$ (4)
In the following, we will call a (separable) $p$-adic Banach space $(X,\|\cdot\|$) admitting a normal basis (equivalently, such that $\|X\|=\left|\mathbb{Q}_{p,\mu}\right|$) a normal $p$-adic Banach space.
Remark 2.5.
A less stringent notion of orthogonality in a $p$-adic normed space $(X,\|\cdot\|)$ is given by the so-called $t$-orthogonality [29]. Specifically, we say that two vectors $x,y$ in $X$ are $t$-orthogonal — and we write $x\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}_{t}y$ — if the condition
$\|\alpha x+\beta y\|\geq t\,\max\big\{\|\alpha x\|,\|\beta y\|\big\},\quad t\in(0,1],$ (5)
holds for every $\alpha,\beta\in\mathbb{Q}_{p,\mu}$. Note that $x\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}_{t}y$ iff $y\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}_{t}x$. In the case where $t=1$, relation (5) entails the orthogonality condition (2); that is, $1$-orthogonality coincides with the — previously introduced — notion of orthogonality in Definition 2.1. We say that $\{x,y\}\subset X\setminus\{0\}$ is a $t$-orthogonal system if $x$ and $y$ are (mutually) $t$-orthogonal vectors, while a (nonempty) subset $\mathfrak{N}$ of $X\setminus\{0\}$ is $t$-orthogonal, if either $\mathfrak{N}$ is a singleton or, for any $x,y\in\mathfrak{N}$, with $x\neq y$, $\{x,y\}$ is a $t$-orthogonal system. Finally, in the same vein of Definition 2.3, a system $\bm{h}\equiv\{h_{i}\}_{i\in I}$ in $X$ is called a $t$-orthogonal basis (of X), if $\bm{h}$ is a $t$-orthogonal system and, additionally, if any element $x$ in $X$ can be expressed, in a unique way, as the sum of a norm-convergent series of the form $x=\sum_{i\in I}\alpha_{i}h_{i}$, where $\alpha_{i}\in\mathbb{Q}_{p,\mu}$.
Remark 2.6.
Let $(X,\|\cdot\|)$ be a $p$-adic Banach space. We recall that the set $c_{0}(I,X)$ — where $I$ is a countable index set, i.e., $I=\{1,\ldots,n\}$ or $I=\mathbb{N}$ — of zero-convergent sequences in $X$ is defined as [7, 22, 23]
$c_{0}(I,X)\coloneqq\{\{x_{i}\}_{i\in I}\mid x_{i}\in X,\;\mbox{$\lim_{i}x_{i}=0$}\}$ (6)
— in the case where $I$ is finite, we set $\lim_{i}x_{i}\equiv 0$. It is possible to prove that $c_{0}(I,X)$, endowed with the so-called sup-norm $\|\cdot\|_{\infty}$, defined for any $x=\{x_{i}\}_{i\in I}\in c_{0}(I,X)$ as
$\|x\|_{\infty}\coloneqq\sup_{i\in I}\|x_{i}\|,$ (7)
is a $p$-adic Banach space [7, 22, 23]. In particular, with $X=\mathbb{Q}_{p,\mu}$, viewed as a one dimensional vector space equipped with the norm $\left|\;\cdot\;\right|$, we have the Banach space
$c_{0}(I,\mathbb{Q}_{p,\mu})\coloneqq\{\{\alpha_{i}\}_{i\in I}\mid\alpha_{i}\in\mathbb{Q}_{p,\mu},\;\mbox{$\lim_{i}\alpha_{i}=0$}\}.$ (8)
We next recall the notion of a $p$-adic Hilbert space [22, 23]. As in the standard complex setting, this notion stems from a suitable definition of an inner product, but, remarkably, in the $p$-adic setting the norm and the inner product are not related in the usual way.
Let $(X,\|\cdot\|)$ be a $p$-adic Banach space (or, more generally, a $p$-adic normed space) over a quadratic extension $\mathbb{Q}_{p,\mu}$ of $\mathbb{Q}_{p}$. We call a map $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle\colon X\times X\rightarrow\mathbb{Q}_{p,\mu}$ a non-Archimedean inner product on $X$, if it is sesquilinear, Hermitian, and satisfies the Cauchy-Schwarz inequality, i.e.,
$\left|\langle x,y\rangle\right|\leq\|x\|\|y\|,\quad\forall x,y\in X.$ (9)
[…]
Definition 2.7.
Let $(X,\|\cdot\|,\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle)$ be an inner-product $p$-adic Banach space. We say that two vectors $x,y$ in $X$ are inner-product-orthogonal (in short, IP-orthogonal), and we write $x\perp y$, if $\langle x,y\rangle=0$.
Remark 2.8.
It is worth observing that in an inner-product $p$-adic Banach space, in general, $\|x\|\neq\sqrt{\left|\langle x,x\rangle\right|}$, i.e., the ultrametric norm does not stem directly from the non-Archimedean inner-product [22, 23]. Accordingly, in the $p$-adic setting, we have two distinct natural notions of orthogonality: the IP-orthogonality and the (previously introduced) norm-orthogonality.
A countable set of vectors $\Psi\equiv\{\psi_{i}\}_{i\in I}$ in an inner product $p$-adic Banach space $(X,\|\cdot\|)$ is said to be an orthonormal basis, if $\Psi$ is a normal basis in $X$ (see Definition 2.3) and if, moreover, $\langle\psi_{i},\psi_{j}\rangle=\delta_{ij}$, for every pair of indices $i,j\in I$. In an inner product $p$-adic Banach space, the existence of an orthonormal basis is not guaranteed. Nevertheless, when such a basis exists, several salient properties of complex Hilbert spaces have analogs in the p-adic setting. It is precisely this observation that led us to set the following [22, 23]:
Definition 2.9.
We call a quadruple $(\mathcal{H},\|\cdot\|,\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle,\Psi)$, where (,,,)($\mathcal{H}$$\|\cdot\|,\langle\hskip 0.9105pt\cdot\hskip 1.70717pt$$\cdot\hskip 0.9105pt\rangle$) is an inner-product $p$-adic Banach space, and $\Psi\equiv\{\psi_{i}\}_{i\in I}$ an orthonormal basis in $\mathcal{H}$, a $p$-adic Hilbert space. We put $\dim(\mathcal{H})\coloneqq\mathrm{card}(I)$.
Remark 2.10.
Let $(\mathcal{H},\|\cdot\|,\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle,\Psi)$ be a $p$-adic Hilbert space ($\dim(\mathcal{H})=\mathrm{card}(I)$). Owing to the continuity of the inner product w.r.t. both its arguments, it is not difficult to verify that any $x\in\mathcal{H}$ admits a (unique) representation — w.r.t. the orthonormal basis $\Psi$ — as the sum on an unconditional norm convergent series of the form
$x=\sum_{i\in I}\langle\psi_{i},x\rangle\psi_{i}.$ (10)
In particular, from relation (10), one can derive the non-Archimedean Parseval identity [22]:
$\|x\|=\max_{i\in I}\left|\langle\psi_{i},x\rangle\right|.$ (11)
Example 2.11.
Consider the $p$-adic Banach space $(c_{0}(I,\mathbb{Q}_{p,\mu}),\|\cdot\|_{\infty})$ (see Remark 2.6). This space admits a normal basis $\bm{e}=\{e_{i}\}_{i\in I}$ consisting of all the sequences $e_{i}$ in $c_{0}(I,\mathbb{Q}_{p,\mu})$ of the form
$e_{1}=(1,0,0,\ldots),\quad e_{2}=(0,1,0,\ldots),\quad e_{3}=(0,0,1,\ldots),\quad\ldots\quad.$ (12)
We can endow $c_{0}(I,\mathbb{Q}_{p,\mu})$ with the so-called canonical inner product associated with $\bm{e}\equiv\{e_{i}\}_{i\in I}$, i.e., the map defined, for every pair of vectors $x=\sum_{i\in I}x_{i}e_{i}$, and $y=\sum_{j\in I}y_{j}e_{j}$ in $c_{0}(I,\mathbb{Q}_{p,\mu})$, as
$c_{0}(I,\mathbb{Q}_{p,\mu})\times c_{0}(I,\mathbb{Q}_{p,\mu})\ni(x,y)\mapsto\langle x,y\rangle_{c}\coloneqq\sum_{i\in I}\overline{x}_{i}y_{i}\in\mathbb{Q}_{p,\mu}.$ (13)
By construction, we have: $\langle e_{i},e_{j}\rangle_{c}=\delta_{ij}$, $\forall i,j\in I$, i.e., $\bm{e}\equiv\{e_{i}\}_{i\in I}$ is an orthonormal basis. Therefore, $(c_{0}(I,\mathbb{Q}_{p,\mu}),\|\cdot\|_{\infty},\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{c},\bm{e})$ is a $p$-adic Hilbert space; in the literature [22, 23, 34, 35] it is sometimes referred to as the coordinate $p$-adic Hilbert space, and denoted by $\mathbb{H}(I)$.
Remark 2.12.
If $\mathcal{H}$ is a $p$-adic Hilbert space, then $\|\mathcal{H}\|=\left|\mathbb{Q}_{p,\mu}\right|$, i.e., $\mathcal{H}$ is a normal $p$-adic Banach space. Conversely, given a normal $p$-adic Banach space $X$, and a normal basis $\bm{e}\equiv\{e_{i}\}_{i\in I}$, we can always endow $X$ with a suitable inner product — the canonical inner product associated with $\{e_{i}\}_{i\in I}$, see Example 2.11 — such that $\bm{e}$ becomes an orthonormal basis, and $X$ a $p$-adic Hilbert space [22, 23].
[…]
  1. (A1)
    $\|Wx\|_{\mbox{\tiny$\mathcal{K}$}}=\|x\|_{\mbox{\tiny$\mathcal{H}$}},\quad\forall x\in\mathcal{H}$;
  2. (A2)
    $\langle Wx,Wy\rangle_{\mbox{\tiny$\mathcal{K}$}}=\langle x,y\rangle_{\mbox{\tiny$\mathcal{H}$}},\quad\forall x,y\in\mathcal{H}$;
i.e., $W$ is an inner-product-preserving surjective isometry [22, 23]. In complete analogy with the standard complex setting, it is possible to prove that two $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$ are isomorphic iff $\dim(\mathcal{H})=\dim(\mathcal{K})$ [22, 23].
Example 2.13.
Consider the coordinate $p$-adic Hilbert space $\mathbb{H}(I)$ introduced in Example 2.11. If $\mathcal{H}$ is a $p$-adic Hilbert space, and $\Psi=\{\psi_{i}\}_{i\in I}$ an orthonormal basis, it is not difficult to see that the map defined as
$W_{\Psi}\colon\mathcal{H}\ni x=\sum_{i\in I}\langle\psi_{i},x\rangle\psi_{i}\mapsto W_{\Psi}(x)\coloneqq\{\langle\psi_{i},x\rangle\}_{i\in I}\equiv\breve{x}\in\mathbb{H}(I)$ (14)
provides an isomorphism (the so-called canonical isomorphism) of $\mathcal{H}$ onto $\mathbb{H}(I)$ [22, 23].
Remark 2.14.
Recall that the topological dual $\mathcal{H}^{\prime}$ of a Hilbert space $\mathcal{H}$ is the set of all bounded linear functionals on $\mathcal{H}$. If $\mathcal{H}$ is a $p$-adic Hilbert space such that $\dim(\mathcal{H})=\infty$, $\mathcal{H}^{\prime}$ is not a $p$-adic Hilbert space, but a $p$-adic Banach space isomorphic to the set $\ell^{\infty}(I,\mathbb{Q}_{p,\mu})$ of bounded sequences in $\mathbb{Q}_{p,\mu}$ [22, 23],
$\ell^{\infty}(I,\mathbb{Q}_{p,\mu})\coloneqq\big\{\eta\equiv\{\eta_{i}\}_{i\in I}\mid\eta_{i}\in\mathbb{Q}_{p,\mu},\,\eta\,\mbox{bounded sequence in $\mathbb{Q}_{p,\mu}$}\big\}.$ (15)
In fact, the mapping defined by
$\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}\colon\mathcal{H}\ni\psi\mapsto\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}\psi\coloneqq\langle\psi,\cdot\hskip 1.13809pt\rangle\in\mathcal{H}^{\prime},$ (16)
provides a conjugate-linear isometry of $\mathcal{H}$ into its dual $\mathcal{H}^{\prime}$, that is surjective iff $\dim(\mathcal{H})<\infty$. The isomorphism between $\mathcal{H}^{\prime}$ and the ultrametric Banach space $\ell^{\infty}(I,\mathbb{Q}_{p,\mu})$ — with $\mathrm{card}(I)=\dim(\mathcal{H}^{\prime})$ — is realized via the surjective isometry defined, for any given orthonormal basis $\Psi\equiv\{\psi_{i}\}_{i\in I}$ in $\mathcal{H}$, by
$\mathcal{L}_{\Psi}\colon\ell^{\infty}(I,\mathbb{Q}_{p,\mu})\ni\xi=\{\xi_{i}\}_{i\in I}\mapsto\sum_{i\in I}\xi_{i}\langle\psi_{i},\cdot\hskip 1.13809pt\rangle=\sum_{i\in I}\xi_{i}\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}\psi_{i}\in\mathcal{H}^{\prime},$ (17)
where, if $I=\mathbb{N}$, the series converges in the weak-topology (see Proposition $3.36$ in [22]).
2.2 Linear operators between $p$-adic Hilbert spaces
We now turn our attention to some classes of bounded operators from a $p$-adic Hilbert space $\mathcal{H}$ into another $p$-adic Hilbert space $\mathcal{K}$. The case where $\mathcal{H}=\mathcal{K}$ has already been analyzed in [22], and the general case we are considering here can be treated by means of a straightforward modification of the methods developed therein.
Hereafter, for the sake of definiteness, it is worth assuming that $\dim(\mathcal{H})=\dim(\mathcal{K})=\infty$, and, accordingly, we will identify the index set $I$ of the previous subsection with $\mathbb{N}$. In all other cases — $\dim(\mathcal{H})<\dim(\mathcal{K})=\infty$, etc. — analogous results hold with rather obvious modifications. For the sake of notational simplicity, we will adopt the shorthand notations $c_{0}$, $\ell^{\infty}$ and $\mathbb{H}$ for $c_{0}(\mathbb{N},\mathbb{Q}_{p,\mu})$, $\ell^{\infty}(\mathbb{N},\mathbb{Q}_{p,\mu})$ and $\mathbb{H}(\mathbb{N})$, respectively.
2.2.1 Bounded and adjointable operators
As in the standard complex setting, linear operators between p-adic Hilbert spaces are also characterized by their domain — a linear space that may not coincide with the original Hilbert space, though it is a linear subspace thereof — which is typically specified prior to defining the operator itself. Let $\mathcal{H}$ and $\mathcal{K}$ be two $p$-adic Hilbert spaces, and let $B\colon\mathrm{dom}(B)\rightarrow\mathcal{K}$, with $\mathrm{dom}(B)\subset\mathcal{H}$, be a linear operator. It is often useful to use the shorthand notation $B\colon\mathcal{H}\dashrightarrow\mathcal{K}$ to highlight the initial space $\mathcal{H}$ and the final space $\mathcal{K}$, rather than the domain $\mathrm{dom}(B)\subset\mathcal{H}$ and range $\mathrm{ran}(B)\coloneqq B(\mathrm{dom}(B))\subset\mathcal{K}$.
We say that $B\colon\mathcal{H}\dashrightarrow\mathcal{K}$ is a bounded operator if
$\|B\|\coloneqq\sup_{\mathrm{dom}(B)\ni x\neq 0}\frac{\|Bx\|}{\|x\|}<\infty.$ (18)
As in the complex setting, $B\colon\mathcal{H}\dashrightarrow\mathcal{K}$ is bounded iff it is continuous [22, 23]; moreover, if $B$ is bounded, it can be extended by continuity to the norm-closure of its domain $\mathrm{dom}(B)$, which is a $p$-adic Banach space, but, in general, not a $p$-adic Hilbert space. Therefore, in the $p$-adic setting (bounded) operators whose domain coincides with the full initial space deserve special attention, and we adopt a ad hoc terminology and notation.
Consequently, we will say that a linear operator $B\colon\mathcal{H}\dashrightarrow\mathcal{K}$ is all-over, if $\mathrm{dom}(B)=\mathcal{H}$, and, in such a case, we will write $B\colon\mathcal{H}\rightarrow\mathcal{K}$; otherwise stated, by writing $B\colon\mathcal{H}\rightarrow\mathcal{K}$ we mean precisely that $\mathrm{dom}(B)=\mathcal{H}$.
We will denote by $\mathcal{B}(\mathcal{H},\mathcal{K})$ the set of all-over, bounded — equivalently, continuous — linear operators from $\mathcal{H}$ to $\mathcal{K}$; i.e., $\mathcal{B}(\mathcal{H},\mathcal{K})\coloneqq\{B\colon\mathcal{H}\rightarrow\mathcal{K}\mid\mbox{$B$ bounded}\}$.
It turns out that, in the $p$-adic setting, the linear operators in $\mathcal{B}(\mathcal{H},\mathcal{K})$ are best described and characterized as matrix operators w.r.t. any given pair of orthonormal bases in the initial space $\mathcal{H}$ and in the final space $\mathcal{K}$. We then briefly introduce this (larger) class of linear operators.
Let $(A_{mn})$, $m,n\in\mathbb{N}$, be an infinite matrix in $\mathbb{Q}_{p,\mu}$ (we will denote by $\mathsf{M}_{\infty}(\mathbb{Q}_{p,\mu})$ the set of all such matrices), and let $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ be orthonormal bases in $\mathcal{H}$ and $\mathcal{K}$, respectively. The matrix $(A_{mn})$ and the orthonormal bases $\Phi$, $\Psi$, determine a matrix operator $\mathrm{op}_{\Phi,\Psi}(A_{mn})\colon\mathcal{H}\dashrightarrow\mathcal{K}$ defined by
$\mathrm{dom}(\mathrm{op}_{\Phi,\Psi}(A_{mn}))\coloneqq$ $\Big\{\mbox{$\zeta=\sum_{n}z_{n}\phi_{n}\in\mathcal{H}\,\,\Big\lvert\,\,\sum_{n}A_{mn}z_{n}$ converges, $\forall m\in\mathbb{N}$, and}$
$\mbox{$\sum_{m}\big(\sum_{n}A_{mn}z_{n}\big)\psi_{m}$ converges too, i.e., $\big\{\sum_{n}A_{mn}z_{n}\big\}_{m\in\mathbb{N}}\in c_{0}$}\Big\},$ (19)
$\mathrm{op}_{\Phi,\Psi}(A_{mn})\zeta\coloneqq\mbox{$\sum_{m}\big(\sum_{n}A_{mn}\langle\phi_{n},\zeta\rangle\big)\psi_{m}$},\quad\forall\zeta\in\mathrm{dom}(\mathrm{op}_{\Phi,\Psi}(A_{mn})).$ (20)
Remark 2.15.
We stress that the matrix operator $\mathrm{op}_{\Phi,\Psi}(A_{mn})\colon\mathcal{H}\dashrightarrow\mathcal{K}$ is — in general — unbounded.
Let us focus on the case of a matrix operator $B=\mathrm{op}_{\Phi,\Psi}(B_{mn})\colon\mathcal{H}\rightarrow\mathcal{K}$, i.e., $\mathrm{dom}(B)=\mathcal{H}$. It is natural to wonder whether such an operator is bounded and, in the case where the answer is affirmative, whether every bounded operator $B\in\mathcal{B}(\mathcal{H},\mathcal{K})$ is a matrix operator of this form.
In analogy with the notation introduced in [22] — where matrix operators acting on a single Hilbert space are considered — we denote by $(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$ the set of the all-over matrix operators from $\mathcal{H}$ to $\mathcal{K}$, associated with the pair of orthonormal bases $\Phi\subset\mathcal{H}$ and $\Psi\subset\mathcal{K}$; i.e., we put
$(\mathcal{H},\mathcal{K})_{\Phi,\Psi}\coloneqq\big\{\mathrm{op}_{\Phi,\Psi}(B_{mn})\colon\mathcal{H}\dashrightarrow\mathcal{K}\mid(B_{mn})\in\mathsf{M}_{\infty}(\mathbb{Q}_{p,\mu}),\,\,\mathrm{dom}(\mathrm{op}_{\Phi,\Psi}(B_{mn}))=\mathcal{H}\big\}.$ (21)
Actually, the set $(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$ does not depend on a particular choice of the pair of orthonormal bases $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}\subset\mathcal{H}$ and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}\subset\mathcal{K}$. Moreover, it can be proven [22] that the matrix operator $\mathrm{op}_{\Phi,\Psi}(B_{mn})$ is all-over iff
$\lim_{m}B_{mn}=0,\;\forall n\in\mathbb{N},\;\;\mbox{and}\;\;\sup_{m,n}\left|B_{mn}\right|<\infty.$ (22)
Consider now a bounded operator $B\in\mathcal{B}(\mathcal{H},\mathcal{K})$. Clearly, we have:
$\lim_{m}\langle\psi_{m},B\phi_{n}\rangle=0,\;\forall n\in\mathbb{N},\,\,\mbox{and}\,\,\sup_{m,n}\left|\langle\psi_{m},B\phi_{n}\rangle\right|\leq\|B\|.$ (23)
Therefore, setting $B_{mn}\equiv\langle\psi_{m},B\phi_{n}\rangle$, the matrix operator $\mathrm{op}_{\Phi,\Psi}(B_{mn})$ is all-over, and, in fact, one can easily check that
$B=\sum_{m}\sum_{n}B_{mn}\langle\phi_{n},\cdot\rangle\psi_{m}=\sum_{n}\sum_{m}B_{mn}\langle\phi_{n},\cdot\hskip 1.13809pt\rangle\psi_{m}=\mathrm{op}_{\Phi,\Psi}(B_{mn}).$ (24)
Here, $(\langle\phi_{n},\cdot\rangle\psi_{m})\zeta\coloneqq\langle\phi_{n},\zeta\rangle\psi_{m}$ — i.e., $\langle\phi_{n},\cdot\rangle\psi_{m}\equiv|\psi_{m}\rangle\langle\phi_{n}|$ in Dirac’s notation — and the iterated series converge w.r.t. the strong operator topology of $\mathcal{B}(\mathcal{H},\mathcal{K})$. Hence, we have that $\mathcal{B}(\mathcal{H},\mathcal{K})\subset(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$.
The crucial observation, at this point, is that the reverse inclusion holds too; i.e., one can prove the following result, which is a natural generalization of Theorem $4.2$ in [22]:
Theorem 2.16.
Let $\mathcal{H}$ and $\mathcal{K}$ be $p$-adic Hilbert spaces. For every pair of orthonormal bases $\Phi=\{\phi_{n}\}_{n\in\mathbb{N}}\subset\mathcal{H}$ and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}\subset\mathcal{K}$, the following relation holds:
$\mathcal{B}(\mathcal{H},\mathcal{K})=\big\{\mathrm{op}_{\Phi,\Psi}(B_{mn})\colon\mathcal{H}\dashrightarrow\mathcal{K}\mid\mbox{$(B_{mn})\in\mathsf{M}_{\infty}(\mathbb{Q}_{p,\mu})$, $\mathrm{dom}(\mathrm{op}_{\Phi,\Psi}(B_{mn}))=\mathcal{H}$}\big\}=(\mathcal{H},\mathcal{K})_{\Phi,\Psi}.$ (25)
Moreover, every bounded operator $B\in\mathcal{B}(\mathcal{H},\mathcal{K})$ can be expressed as in (24), and we have:
$\|B\|=\sup_{m,n}\left|B_{mn}\right|=\sup_{n}\|B\phi_{n}\|.$ (26)
Proof.
See Appendix A. ∎
By Theorem 2.16, $\mathcal{B}(\mathcal{H},\mathcal{K})$ coincides with the set $(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$ of the all-over matrix operators from $\mathcal{H}$ into $\mathcal{K}$ — independently of the choice of the orthonormal bases $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ — and, moreover, the infinite matrices in $\mathsf{M}_{\infty}(\mathbb{Q}_{p,\mu})$ corresponding to bounded operators in $\mathcal{B}(\mathcal{H},\mathcal{K})$ admit a complete characterization.
In the standard complex setting, every bounded operator admits a (Hilbert space) adjoint, which is also a bounded operator. In the $p$-adic setting — since the dual $\mathcal{H}^{\prime}$ of $\mathcal{H}$ cannot be identified with $\mathcal{H}$ itself — the notion of operator adjoint is not as straightforward as in the complex case. However, there is a distinguished class of bounded operators — called adjointable (bounded) operators — that may be regarded as the true $p$-adic counterpart of the bounded linear operators between complex Hilbert spaces.
Let us first recall that a bounded operator $B\in\mathcal{B}(\mathcal{H},\mathcal{K})$ will admit a (generalized) adjoint $B^{\prime}\in\mathcal{B}(\mathcal{K}^{\prime},\mathcal{H}^{\prime})$ defined by
$\mathcal{K}^{\prime}\ni\omega\mapsto B^{\prime}\omega\coloneqq\omega\circ B\in\mathcal{H}^{\prime}.$ (27)
However, it is possible to single out, within $\mathcal{B}(\mathcal{H},\mathcal{K})$, a suitable class of bounded operators that admit a proper Hilbert space adjoint; i.e., a suitable bounded operator in $\mathcal{B}(\mathcal{K},\mathcal{H})$. In fact, with every $B\in\mathcal{B}(\mathcal{H},\mathcal{K})$ we can associate a linear operator $B^{\dagger}\colon\mathcal{K}\dashrightarrow\mathcal{H}$ — the so-called pseudo-adjoint of $B$ — defined as follows. We first put
$\mathrm{dom}\big(B^{\dagger}\big)\coloneqq\{\psi\in\mathcal{K}\mid\mbox{there exists $\zeta\in\mathcal{H}$ s.t.\ $\langle\psi,B\phi\rangle=\langle\zeta,\phi\rangle$, $\forall\phi\in\mathcal{H}=\mathrm{dom}(B)$}\}.$ (28)
It is clear that $\mathrm{dom}(B^{\dagger})$ is a linear subspace of $\mathcal{K}$; it can be regarded — via the conjugate-linear isometry $\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}$ (see Remark 2.14) — as the set of all vectors $\psi\in\mathcal{K}$ such that the bounded functional $\langle\psi,B(\cdot)\rangle\in\mathcal{H}^{\prime}$ can be identified with an element $\eta$ of $\mathcal{H}$, i.e.,
$\psi\in\mathrm{dom}(B^{\dagger})\;\;\iff\;\;\langle\psi,B(\cdot)\rangle=\langle\eta,\cdot\hskip 1.13809pt\rangle=\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}\eta.$ (29)
The condition
$\big(\mathcal{J}_{\mbox{\tiny$\mathcal{K}$}}\psi\big)(B\phi)=\langle\psi,B\phi\rangle=\langle\eta,\phi\rangle=\big(\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}\eta\big)(\phi),\quad\forall\psi\in\mathrm{dom}(B^{\dagger}),\,\forall\phi\in\mathcal{H}$ (30)
uniquely determines a linear operator $B^{\dagger}\colon\mathcal{K}\dashrightarrow\mathcal{H}$, whose domain $\mathrm{dom}\big(B^{\dagger}\big)$ is defined by (28), by setting $B^{\dagger}\psi\coloneqq\eta$ (in fact, since $\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}$ is a conjugate linear isometry, there exists a unique vector $\eta\in\mathcal{H}$ such that $\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}\eta=\big(\mathcal{J}_{\mbox{\tiny$\mathcal{K}$}}\psi\big)\circ B$). Moreover, by (30), one has that
$\|B^{\dagger}\psi\|=\|\eta\|=\|\mathcal{J}_{\mbox{\tiny$\mathcal{H}$}}\eta\|=\|(\mathcal{J}_{\mbox{\tiny$\mathcal{K}$}}\psi)\circ B\|\leq\|\mathcal{J}_{\mbox{\tiny$\mathcal{K}$}}\psi\|\,\|B\|=\|\psi\|\,\|B\|,$ (31)
i.e., $\|B^{\dagger}\|\leq\|B\|$ and we see that $B^{\dagger}$ is a bounded operator.
One can show (along the lines traced in [22]) that either $\mathrm{dom}\big(B^{\dagger}\big)=\mathcal{K}$ or $\mathrm{dom}\big(B^{\dagger}\big)$ is a proper closed subspace of $\mathcal{K}$. Taking into account this fact, we now set the following:
Definition 2.17.
Let $B\in\mathcal{B}(\mathcal{H},\mathcal{K})$. We say that $B$ is adjointable if $\mathrm{dom}(B^{\dagger})=\mathcal{K}$, i.e., if the bounded operator $B^{\dagger}$ — the pseudo-adjoint of $B$ — is all-over. In such a case, we write $B^{\ast}$ instead of $B^{\dagger}$ and call $B^{\ast}$ the adjoint of $B$ in $\mathcal{B}(\mathcal{H},\mathcal{K})$.
In what follows, we will denote by $\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$ the set of all adjointable bounded operators from $\mathcal{H}$ to $\mathcal{K}$. We now aim to provide a suitable matrix characterization of an adjointable operator. In particular, we prove the following:
Theorem 2.18.
Let $B\in\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$, and let $B^{\ast}\colon\mathcal{K}\rightarrow\mathcal{H}$ be its adjoint. Then, $B^{\ast}\in\mathcal{B}_{\mathrm{ad}}(\mathcal{K},\mathcal{H})$; i.e., $B^{\ast}$ is both bounded and adjointable. Specifically, given any pair of orthonormal bases $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$ and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{K}$, if $B=\mathrm{op}_{\Phi,\Psi}(B_{mn})\in\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$, then
$B^{\ast}=\mathrm{op}_{\Psi,\Phi}(B^{\ast}_{mn}),$ (32)
where we have put $B_{mn}^{\ast}=\overline{B_{nm}}$, and the coefficients of the matrix $(B_{mn})$ associated with $B$ satisfy the following conditions:
  1. (1)
    $\sup_{m,n}\left|B_{mn}\right|<\infty$;
  2. (2)
    $\lim_{m}B_{mn}=0$, $\forall n\in\mathbb{N}$;
  3. (3)
    $\lim_{n}B_{mn}=0$, $\forall m\in\mathbb{N}$.
Moreover, it is clear that
$\|B^{\ast}\|=\sup_{m,n}\left|B_{mn}\right|=\|B\|.$ (33)
Conversely, if $B\colon\mathcal{H}\dashrightarrow\mathcal{K}$ is a matrix operator — i.e., if $B=\mathrm{op}_{\Phi,\Psi}(B_{mn})$, for some pair of orthonormal bases $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}\subset\mathcal{H}$ and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}\subset\mathcal{K}$ — and, moreover, the coefficients of the matrix $(B_{mn})\in\mathsf{M}_{\infty}(\mathbb{Q}_{p,\mu})$ satisfy conditions (1)(3) above, then we have that $B\colon\mathcal{H}\rightarrow\mathcal{K}\in\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$, and, therefore, the adjoint $B^{\ast}\colon\mathcal{K}\rightarrow\mathcal{H}\in\mathcal{B}_{\mathrm{ad}}(\mathcal{K},\mathcal{H})$ of $B$ can be expressed as a matrix operator of the form $B^{\ast}=\mathrm{op}_{\Psi,\Phi}(B^{\ast}_{mn})=\mathrm{op}_{\Psi,\Phi}(\overline{B_{nm}})$.
Proof.
See Appendix A. ∎
Remark 2.19.
From Theorem 2.18 it follows that, if $B=\mathrm{op}_{\Phi,\Psi}(B_{mn})$ is an adjointable operator, then $B^{\ast}$ is adjointable too, and $(B^{\ast})^{\ast}=B$. Moreover, $\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$ is a linear subspace of $\mathcal{B}(\mathcal{H},\mathcal{K})$, as one can easily check by observing that for all $A,B\in\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$, and all $\alpha\in\mathbb{Q}_{p,\mu}$, $\alpha A$ and $A+B$ are in $\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$.
2.2.2 The $p$-adic trace/Hilbert-Schmidt class
As is well known, in the standard complex setting, the trace class and the Hilbert-Schmidt class, are two distinct operator spaces: a Banach space, in the first case, and a Hilbert space (endowed with the Hilbert-Schmidt scalar product) in the second [28], the former space being strictly contained in the latter in the case where the carrier Hilbert space is infinite-dimensional. In the $p$-adic setting, instead, we have a unique operator space that plays the role of both the trace and the Hilbert-Schmidt class. However, we will first introduce this space as a $p$-adic Banach space — the so-called $p$-adic trace class $\mathcal{T}(\mathcal{H},\mathcal{K})$ — and we will next argue that this space, endowed with a suitable scalar product, becomes a $p$-adic Hilbert space; i.e., the $p$-adic Hilbert-Schmidt class. Another distinguishing feature of the $p$-adic trace class $\mathcal{T}(\mathcal{H},\mathcal{K})$ is that it is a Banach space w.r.t. the standard operator norm (rather than some ‘trace norm’).
Let $\mathcal{H}$ and $\mathcal{K}$ be $p$-adic Hilbert spaces, with orthonormal bases $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$, and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$, respectively. Extending, in a natural way, the definition [22, 23] of the trace class $\mathcal{T}(\mathcal{H})$ of a single $p$-adic Hilbert space, we now introduce the set of trace class operators from $\mathcal{H}$ to $\mathcal{K}$ as follows:
$\mathcal{T}(\mathcal{H},\mathcal{K})\coloneqq\mbox{$\{\mathrm{op}_{\Phi,\Psi}(T_{mn})\colon\mathcal{H}\dashrightarrow\mathcal{K}\mid T_{mn}\in\mathsf{M}_{\infty}(\mathbb{Q}_{p,\mu})\,\,\text{s.t.}\,\,\lim_{m+n}T_{mn}=0\}$}.$ (34)
At this point, a further comment is in order. The reader will be familiar with the standard definition of a trace class operator on a complex Hilbert space, which crucially relies on the notion of positive operator and, accordingly, of the absolute value of a bounded operator (see, e.g., Chapter 3 of [28]). In the $p$-adic setting there is no meaningful notion of positivity, and the previous definition of a trace class operator turns out to be the most convenient one.
Remark 2.20.
Note that the defining condition for a trace class operator in the $p$-adic setting — i.e., $T=\mathrm{op}_{\Phi,\Psi}(T_{mn})\in\mathcal{T}(\mathcal{H},\mathcal{K})$ iff $\lim_{m+n}T_{mn}=0$ — involves a pair of orthonormal bases $\Phi$, $\Psi$, because $T$ is defined as a matrix operator, but the property of being a trace class operator is actually independent of the choice of this pair of bases; see below. Also note that, if $\lim_{m+n}T_{mn}=0$, then the sum $\sum_{m}T_{mm}$ of all ‘diagonal entries’ is a convergent series in $\mathbb{Q}_{p,\mu}$.
Exploiting the defining condition (34), and by suitably adapting to the present setting the line of reasoning in Sec. VI of [22] (see, in particular, Propositions $6.2$ and $6.7$, and Corollary $6.8$), the forthcoming result can be proved. We will denote by $\mathcal{V}$ any third Hilbert space over $\mathbb{Q}_{p,\mu}$.
Theorem 2.21.
The definition (34) of the set $\mathcal{T}(\mathcal{H},\mathcal{K})$ does not depend on the choice of the pair of orthonormal bases $\Phi$, $\Psi$. Moreover, $\mathcal{T}(\mathcal{H},\mathcal{K})$ is a Banach subspace of $\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$, and
$T\in\mathcal{T}(\mathcal{H},\mathcal{K})\subset\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$ $\implies T^{\ast}\in\mathcal{T}({\mathcal{K}},{\mathcal{H}}),$
$T\in\mathcal{T}(\mathcal{H},\mathcal{K}),\ A\in\mathcal{B}_{\mathrm{ad}}(\mathcal{K},\mathcal{V}),\ B\in\mathcal{B}_{\mathrm{ad}}(\mathcal{V},\mathcal{H})$ $\implies AT\in\mathcal{T}({\mathcal{H}},{\mathcal{V}}),\ TB\in\mathcal{T}({\mathcal{V}},{\mathcal{K}}).$ (35)
Finally, every trace class operator $T=\mathrm{op}_{\Phi,\Psi}(T_{mn})\in\mathcal{T}(\mathcal{H},\mathcal{K})$ is traceable; that is, for any pair of orthonormal bases $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{K}$, the trace $\mathrm{tr}(T)$ of $T$ — defined as the sum of the convergent series
$\mathrm{tr}(T)\coloneqq\sum_{m}\langle\psi_{m},T\phi_{m}\rangle=\sum_{m}T_{mm}\in\mathbb{Q}_{p,\mu},$ (36)
— does not depend on the choice of the orthonormal bases $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$.
Remark 2.22.
Let $\mathrm{tr}(\hskip 0.56905pt\cdot\hskip 0.56905pt)\colon\mathcal{T}(\mathcal{H},\mathcal{K})\rightarrow\mathbb{Q}_{p,\mu}$ be the map that associates, with every operator $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$, its trace $\mathrm{tr}(T)$ as defined in (36). One can easily check that $\mathrm{tr}(\hskip 0.56905pt\cdot\hskip 0.56905pt)$ is a linear functional on $\mathcal{T}(\mathcal{H},\mathcal{K})$; namely, for any $S,T\in\mathcal{T}(\mathcal{H},\mathcal{K})$ and $\alpha\in\mathbb{Q}_{p,\mu}$,
$\mathrm{tr}(S+T)=\mathrm{tr}(S)+\mathrm{tr}(T),\quad\mathrm{tr}(\alpha T)=\alpha\mathrm{tr}(T).$ (37)
Moreover, for every $T\colon\mathcal{H}\rightarrow\mathcal{K}\in\mathcal{T}(\mathcal{H},\mathcal{K})$, the adjoint $T^{\ast}\in\mathcal{T}(\mathcal{K},\mathcal{H})$ satisfies the usual relation $\mathrm{tr}(T^{\ast})=\overline{\mathrm{tr}(T)}$.
Exactly as in the standard complex setting, a linear operator $C\colon\mathcal{H}\rightarrow\mathcal{K}$ is said to be compact if it maps the unit ball $\mathcal{H}_{1}\coloneqq\{\psi\in\mathcal{H}\mid\|\psi\|\leq 1\}$ of $\mathcal{H}$ onto a precompact subset $C(\mathcal{H}_{1})$ of $\mathcal{K}$; that is, if $\overline{C(\mathcal{H}_{1})}^{\|\cdot\|}$ is a compact subset of $\mathcal{K}$ [22, 27]. In the following, we will denote by $\mathcal{C}(\mathcal{H},\mathcal{K})\subset\mathcal{B}(\mathcal{H},\mathcal{K})$ the set of all compact (bounded) operators from $\mathcal{H}$ to $\mathcal{K}$.
A special class of compact operators is formed by the so-called block-finite operators [22]. Precisely, a matrix operator $B=\mathrm{op}_{\Phi,\Psi}(B_{mn})\colon\mathcal{H}\dashrightarrow\mathcal{K}$ is said to be block-finite w.r.t. a certain pair of orthonormal bases $\Phi\subset\mathcal{H}$, $\Psi\subset\mathcal{K}$ — and we denote the set of all such operators by $\mathscr{B}_{\Phi,\Psi}$ — if there exists some $k\in\mathbb{N}$ such that
$\max\{m,n\}>k\;\;\implies\;\;B_{mn}=0.$ (38)
A block-finite operator has finite rank; therefore, it is adjointable and of trace class.
Theorem 2.23.
The following characterization of the trace class $\mathcal{T}(\mathcal{H},\mathcal{K})$ holds true:
$\mathcal{T}(\mathcal{H},\mathcal{K})=\mathcal{C}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})\coloneqq\mathcal{C}(\mathcal{H},\mathcal{K})\cap\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K}).$ (39)
Moreover, for every pair of orthonormal bases $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{K}$, $\mathcal{T}(\mathcal{H},\mathcal{K})$ coincides with the uniform closure of the set $\mathscr{B}_{\Phi,\Psi}$ of block-finite operators w.r.r.t the orthonormal bases $\Phi$, $\Psi$; i.e.,
$\mathcal{T}(\mathcal{H},\mathcal{K})=\overline{\mathscr{B}_{\Phi,\Psi}}^{\|\cdot\|}.$ (40)
Proof.
The first claim is proved likewise Theorem $6.35$ in [22], whereas the second is a straightforward adaptation of Theorem $6.34$ in [22] to our specific case. ∎
As anticipated, in the concluding part of this subsection, we argue that $\mathcal{T}(\mathcal{H},\mathcal{K})$, endowed with a suitable sesquilinear form, has a natural structure of a $p$-adic Hilbert space, namely, the $p$-adic Hilbert-Schmidt space. In fact, let us introduce the sesquilinear form
$\mathcal{T}(\mathcal{H},\mathcal{K})\times\mathcal{T}(\mathcal{H},\mathcal{K})\ni(S,T)\rightarrow\langle S,T\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}\coloneqq\mathrm{tr}(S^{\ast}T)\in\mathbb{Q}_{p,\mu}.$ (41)
Here, note that, by relations (35), if $S,T\in\mathcal{T}(\mathcal{H},\mathcal{K})$, then $S^{\ast}T\colon\mathcal{H}\rightarrow\mathcal{H}\in\mathcal{T}(\mathcal{H})\equiv\mathcal{T}(\mathcal{H},\mathcal{H})$, so that the preceding mapping is well defined. It is clear that $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}$ is Hermitian, as follows by observing that
$\langle S,T\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\mathrm{tr}(S^{\ast}T)=\overline{\mathrm{tr}(T^{\ast}S)}=\overline{\langle T,S\rangle}_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}.$ (42)
Following the terminology adopted in [22], we call the Hermitian sesquilinear form $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}$ the Hilbert-Schmidt product of $\mathcal{T}(\mathcal{H},\mathcal{K})$.
The final step is the explicit construction of an orthonormal basis for the inner product $p$-adic Banach space $\big(\mathcal{T}(\mathcal{H},\mathcal{K}),\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}\big)$. To this end, given any pair of orthonormal bases $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}\subset\mathcal{H}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}\subset\mathcal{K}$, we introduce the family of matrix operators $\big\{\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\colon\mathcal{H}\rightarrow\mathcal{K}\big\}_{j,k\in\mathbb{N}}$ defined by
$\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\coloneqq\mathrm{op}_{\Phi,\Psi}\big(\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}_{mn}}{\operatorname{\mathnormal{E}}}\big),\quad\mbox{where $\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}_{mn}}{\operatorname{\mathnormal{E}}}=\delta_{jm}\delta_{kn}$};$ (43)
i.e., in Dirac’s notation, we have: $\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}=|\psi_{j}\rangle\langle\phi_{k}|\equiv\langle\phi_{k},\cdot\hskip 1.13809pt\rangle\psi_{j}$. Note that, for every $j,k\in\mathbb{N}$, $\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}$, is a rank-one operator and, therefore, it is in the trace class $\mathcal{T}(\mathcal{H},\mathcal{K})$.
Eventually, we are ready to formulate the following result, which is a natural generalization of Theorem $6.38$ in [22], that holds in the single Hilbert space setting.
Theorem 2.24.
The quadruple
$\Big(\mathcal{T}(\mathcal{H},\mathcal{K}),\|\cdot\|,\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}},\big\{\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\big\}_{j,k\in\mathbb{N}}\Big)$ (44)
— where $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}$ is the Hilbert-Schmidt product defined in (41), and $\big\{\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\big\}_{j,k\in\mathbb{N}}$ the set of trace class operators introduced in (43), that form an orthonormal basis in $\mathcal{T}(\mathcal{H},\mathcal{K})$ — is a $p$-adic Hilbert space.
Proof.
See Appendix A. ∎
3 Construction of the tensor product
In this section, we proceed to the construction of the tensor product of $p$-adic Hilbert spaces. As in the real or complex setting, we first define the algebraic tensor product of two vector spaces over $\mathbb{Q}_{p,\mu}$. We stress that, even if we deal with $p$-adic Hilbert spaces, in this preliminary stage we do not need more than the mere linear properties of these spaces. Next, to endow the algebraic tensor product with a metric structure, we introduce an appropriate ultrametric norm. This is a nontrivial step, because what is the appropriate norm is not clear in advance, and, in fact, it turns out that the analogy with the tensor product of complex Hilbert spaces is not helpful in this case. Finally, after taking the completion w.r.t. the metric induced by this norm, we bring up a suitable inner product and an orthonormal basis, in such a way as to obtain eventually a new $p$-adic Hilbert space; i.e., the tensor product of the initial $p$-adic Hilbert spaces.
Definition 3.1.
Let $X,Y,Z$ be vector spaces over the field $\mathbb{Q}_{p,\mu}$. Recall that a map $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt$ from the Cartesian product $X\times Y$ into $Z$ is said to be bilinear if it is linear w.r.t. each of its arguments; that is
$\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(\alpha_{1}x_{1}+\alpha_{2}x_{2},y)$ $=\alpha_{1}\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x_{1},y)+\alpha_{2}\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x_{2},y),$
$\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x,\beta_{1}y_{1}+\beta_{2}y_{2})$ $=\beta_{1}\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x,y_{1})+\beta_{2}\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x,y_{2}),$ (45)
for all vectors $x,x_{1},x_{2}\in X$, $y,y_{1},y_{2}\in Y$, and all scalars $\alpha_{i},\beta_{i}\in\mathbb{Q}_{p,\mu}$, $i=1,2$. We will denote by $\mathsf{Bil}\hskip 0.85358pt(X\times Y,Z)$ the vector space of all bilinear maps from $X\times Y$ to $Z.$ In the special case where $Z=\mathbb{Q}_{p,\mu}$ (the one-dimensional $\mathbb{Q}_{p,\mu}$-linear space), we put $\mathsf{Bil}\hskip 0.85358pt(X\times Y)\equiv\mathsf{Bil}\hskip 0.85358pt(X\times Y,\mathbb{Q}_{p,\mu})$ (the space of bilinear forms on $X\times Y$).
The algebraic tensor product $X\otimes_{\alpha}Y$ of the vector spaces $X$ and $Y$ can be defined as a space of linear functionals on $\mathsf{Bil}\hskip 0.85358pt(X\times Y)$, in the usual fashion [26].
Definition 3.2.
Let $X,Y$ be vector spaces over $\mathbb{Q}_{p,\mu}$, and let $x\in X$, $y\in Y$. The simple tensor $x\otimes y\colon\mathsf{Bil}\hskip 0.85358pt(X\times Y)\rightarrow\mathbb{Q}_{p,\mu}$ is the linear functional defined by
$(x\otimes y)(\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt)\coloneqq\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x,y),$ (46)
for every bilinear form $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt\in\mathsf{Bil}\hskip 0.85358pt(X\times Y)$.
The algebraic tensor product $X\otimes_{\alpha}Y$ is defined as the linear subspace of the algebraic dual space $\mathsf{Bil}\hskip 0.85358pt(X\times Y)^{\vee}$ spanned by all simple tensors $x\otimes y$. Therefore, a generic element $u\in X\otimes_{\alpha}Y$ can be expressed (not uniquely) as a finite linear combination of the form
$u=\sum_{i=1}^{n}\lambda_{i}\,x_{i}\otimes y_{i}$ (47)
— with $n\in\mathbb{N}$, $\lambda_{i}\in\mathbb{Q}_{p,\mu},x_{i}\in X$ and $y_{i}\in Y$, $i\in\{1,\ldots,n\}$ — so that
$u(\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt)=\sum_{i=1}^{n}\lambda_{i}\,\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x_{i},y_{i})\in\mathbb{Q}_{p,\mu}.$ (48)
Next result is the so-called universal property of the algebraic tensor product [26].
Theorem 3.3.
For every bilinear map $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt:X\times Y\to Z$, there exists a unique linear map $B\colon X\otimes_{\alpha}Y\to Z$ such that $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x,y)=B(x\otimes y)$, for all $x\in X$ and $y\in Y$.
Proof.
Let $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt:X\times Y\to Z$ be a bilinear map and define $B:X\otimes_{\alpha}Y\to Z$ by
$B\Big(\sum_{i=1}^{n}x_{i}\otimes y_{i}\Big)=\sum_{i=1}^{n}\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x_{i},y_{i}).$ (49)
Suppose now that $\sum_{i=1}^{n}x_{i}\otimes y_{i}=0.$ Then, for each linear functional $\varphi$ on $Z,$ the composition $\varphi\circ B$ is a bilinear functional on $X\times Y$ and
$\varphi\Big(\sum_{i=1}^{n}\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x_{i},y_{i})\Big)=\sum_{i=1}^{n}\varphi\circ\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt(x_{i},y_{i})=0.$ (50)
Therefore, $B$ is well defined and the bilinear map $\hskip 0.56905pt\mathfrak{B}\hskip 0.56905pt$ associated with the linear map $B$ yields an identification of $\mathsf{Bil}\hskip 0.85358pt(X\times Y,Z)$ and linear maps from $X\otimes_{\alpha}Y$ to Z. ∎
We now list some useful properties of the tensor product.
Proposition 3.4.
The mapping $X\times Y\ni(x,y)\mapsto x\otimes y\in X\otimes_{\alpha}Y$ satisfies the following properties:
  1. (i)
    $(x_{1}+x_{2})\otimes y=x_{1}\otimes y+x_{2}\otimes y$;
  2. (ii)
    $x\otimes(y_{1}+y_{2})=x\otimes y_{1}+x\otimes y_{2}$;
  3. (iii)
    $\lambda(x\otimes y)=(\lambda x)\otimes y=x\otimes(\lambda y)$;
  4. (iv)
    $0\otimes y=x\otimes 0=0$;
  5. (v)
    if $E\subset X$ and $F\subset Y$ are linearly independent sets, then $\{x\otimes y:x\in E,y\in F\}$ is a linearly independent subset of $X\otimes_{\alpha}Y$;
  6. (vi)
    if $x_{1},\dots,x_{n}\in X$ and $y_{1},\dots,y_{n}\in Y$ — with $y_{1},\dots,y_{n}$ linearly independent and $x_{i}\neq 0$, for all $i\in\{1,\ldots,n\}$ — then $\{x_{1}\otimes y_{1},\dots,x_{n}\otimes y_{n}\}$ is linearly independent set.
Proof.
Properties (i)(iv) hold by basic calculations. The proof of (v) and (vi) follow by a slight adaptation of the analogous results from Proposition 1.1 in [26] and Theorem 10.1.6 in [29]. ∎
By (47) and (iii), every tensor $u\in X\otimes_{\alpha}Y$ can be written in the form
$u=\sum_{i=1}^{n}x_{i}\otimes y_{i},$ (51)
for some $n\in\mathbb{N}$. Moreover, in the case where $X$ and $Y$ are finite-dimensional vector spaces, by (v) we have that
$\dim(X\otimes_{\alpha}Y)=\dim(X)\dim(Y).$ (52)
Remark 3.5.
We stress that, even if decomposition (51) is not unique, there is a smallest number $\mathrm{rk}\hskip 0.85358pt(u)$ — called the rank of the tensor $u\in X\otimes_{\alpha}Y$ — of simple tensors (rank-one tensors) that appear in this decomposition.
In addition to the properties (i)(vi) listed in Proposition 3.4, we have the following further useful fact, namely, a characterization of the zero vector in $X\otimes_{\alpha}Y$.
Proposition 3.6.
Let $X,Y$ be vector spaces over $\mathbb{Q}_{p,\mu}$, and let $X^{\vee},Y^{\vee}$ the associated algebraic dual spaces. Then, the following statements are equivalent for $u=\sum_{i=1}^{n}x_{i}\otimes y_{i}\in X\otimes_{\alpha}Y$:
  1. (vii-1)
    $u=0$;
  2. (vii-2)
    $\sum_{i=1}^{n}\hskip 0.56905pt\mathfrak{X}\hskip 0.56905pt(x_{i})\,\hskip 0.56905pt\mathfrak{Y}\hskip 0.56905pt(y_{i})=0$, for every $\hskip 0.56905pt\mathfrak{X}\hskip 0.56905pt\in X^{\vee}$, $\hskip 0.56905pt\mathfrak{Y}\hskip 0.56905pt\in Y^{\vee}$;
  3. (vii-3)
    $\sum_{i=1}^{n}\hskip 0.56905pt\mathfrak{X}\hskip 0.56905pt(x_{i})\,y_{i}=0$, for every $\hskip 0.56905pt\mathfrak{X}\hskip 0.56905pt\in X^{\vee}$;
  4. (vii-4)
    $\sum_{i=1}^{n}\hskip 0.56905pt\mathfrak{Y}\hskip 0.56905pt(y_{i})\,x_{i}=0$, for every $\hskip 0.56905pt\mathfrak{Y}\hskip 0.56905pt\in Y^{\vee}$.
Proof.
The proof is again a natural adaptation of that in the complex setting, see e.g. Proposition 1.2 in [26]. ∎
[…]
Therefore, the next step of our construction is to endow the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ with a suitable ultrametric norm, in such a way as to obtain, upon completion, a suitable $p$-adic Banach space.
Recall [40] that in the standard setting where the tensor product $\mathscr{H}\otimes\mathscr{K}$ of two complex Hilbert spaces $\mathscr{H}$, $\mathscr{K}$ is considered, the norm of the Hilbert space $\mathscr{H}\otimes\mathscr{K}$ is precisely the one canonically associated with the scalar product of this space. Otherwise stated, a suitable scalar product is defined on the algebraic tensor product $\mathscr{H}\otimes_{\alpha}\mathscr{K}$, so obtaining a pre-Hilbert space, and the (Hilbert) tensor product $\mathscr{H}\otimes\mathscr{K}$ is then achieved as the Hilbert space completion of this space. This approach cannot be pursued in the $p$-adic setting, because, as noted previously, the norm and the scalar product are not related in the usual fashion — but only by the Cauchy-Schwarz inequality — in this case. It is also worth noting that the canonical (Hilbert space) norm on $\mathscr{H}\otimes\mathscr{K}$ does not coincide with the projective norm [26] defined on the algebraic tensor product $\mathscr{H}\otimes_{\alpha}\mathscr{K}$ (the linear spaces $\mathscr{H}$, $\mathscr{K}$ being endowed with their respective Hilbert space norms); namely, the (Hilbert space) norm on $\mathscr{H}\otimes\mathscr{K}$ does not coincide with the following:
$\|u\|_{\pi}\coloneqq\inf\bigg\{\sum_{i=1}^{n}\|x_{i}\|_{\mathscr{H}}\,\|y_{i}\|_{\mathscr{K}}\,\,\bigg\lvert\,\,u=\sum_{i=1}^{n}x_{i}\otimes y_{i},\,x_{i}\in\mathscr{H},y_{i}\in\mathscr{K}\bigg\}.$ (53)
Here, the infimum is taken over all possible representations of $u$.
A natural non-Archimedean counterpart of this norm, compatible with the strong triangle inequality characterizing the $p$-adic setting, is given by the following candidate norm:
$\|u\|_{\pi}\coloneqq\inf\bigg\{\max_{0\leq i\leq n}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}\,\,\bigg\lvert\,\,u=\sum_{i=1}^{n}x_{i}\otimes y_{i},\,x_{i}\in\mathcal{H},y_{i}\in\mathcal{K}\bigg\}.$ (54)
Here, the infimum is taken, once again, over all possible representations of $u\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$.
We now demonstrate that it indeed defines a non-Archimedean norm. To this end, we first prove that the mapping $\mathcal{H}\otimes_{\alpha}\mathcal{K}\ni u\mapsto\|u\|_{\pi}\in\mathbb{R^{+}}$ in (54) is an ultrametric seminorm on the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}.$
Proposition 3.7.
Let $\mathcal{H}$ and $\mathcal{K}$ be two $p$-adic Hilbert spaces. Then, $\|\cdot\|_{\pi}$ is an ultrametric seminorm on the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}.$
Proof.
Starting with linearity, we have to prove that $\|\lambda u\|_{\pi}=\left|\lambda\right|\|u\|_{\pi}$. For $\lambda=0$, this is trivial. If $\lambda\neq 0,$ then $\lambda u=\sum_{i=1}^{n}(\lambda x_{i})\otimes y_{i}$ is a possible representation of $\lambda u$ and, hence, we have:
$\begin{split}\|\lambda u\|_{\pi}\leq\max_{i}\|\lambda x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,&\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}=\max_{i}\left|\lambda\right|\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}=\left|\lambda\right|\max_{i}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}\\ &\implies\|\lambda u\|_{\pi}\leq\left|\lambda\right|\|u\|_{\pi}.\end{split}$ (55)
On the other hand we can write $\|u\|_{\pi}=\|\lambda^{-1}\lambda u\|_{\pi}$ which entails
$\|\lambda^{-1}\lambda u\|_{\pi}\leq\left|\lambda\right|^{-1}\|\lambda u\|_{\pi}\implies\left|\lambda\right|\|u\|_{\pi}\leq\|\lambda u\|_{\pi}.$ (56)
Thus from (55) and (56) we have $\|\lambda u\|_{\pi}=\left|\lambda\right|\|u\|_{\pi}.$ It remains to prove the $\|\cdot\|_{\pi}$ strong triangular inequality. For $\epsilon>0$ and let $u=\sum_{i=1}^{n}x_{i}\otimes y_{i}$ be a possible representation of $u$ and $v=\sum_{i=n+1}^{n+m}x_{i}\otimes y_{i}$ be a possible representation of $v$ such that
$\max_{0\leq i\leq n}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}\leq\|u\|_{\pi}+\epsilon\ \ \ \text{and}\ \ \ \max_{n+1\leq i\leq n+m}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}\leq\|v\|_{\pi}+\epsilon.$ (57)
Observe that $\sum_{i=1}^{n+m}x_{i}\otimes y_{i}$ is a possible representation of $u+v$ and therefore
$\begin{split}\|u+v\|_{\pi}\leq\max_{1\leq i\leq n+m}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}&=\max\{\max_{0\leq i\leq n}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}},\max_{n+1\leq i\leq n+m}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}\}\\ &\leq\max\{\|u\|_{\pi}+\epsilon,\|v\|_{\pi}+\epsilon\}.\end{split}$ (58)
Since this is true for every $\epsilon>0$ we have $\|u+v\|_{\pi}\leq\max\{\|u\|_{\pi},\|v\|_{\pi}\}$ as desired. ∎
Next, to establish that $\|\cdot\|_{\pi}$ is an ultrametric norm, we need to prove that $\|x\|_{\pi}=0$ iff $x=0$. To this end, we start by proving the following fact:
Lemma 3.8.
If $\{x_{i}\}_{i\in I}$ is a finite norm-orthogonal system in $\mathcal{H}$, then
$\bigg\|\sum_{i\in I}x_{i}\otimes y_{i}\bigg\|_{\pi}=\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}},$ (59)
for any choice of the set $\{y_{i}\}_{i\in I}\subset\mathcal{K}.$
Proof.
The proof of this result mimics, to some extent, that of Lemma 10.2.9 in [29], but requires some adaptations to our present case. Inequality $\|\sum_{i\in I}x_{i}\otimes y_{i}\|_{\pi}\leq\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\cdot\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}$ holds by definition of ultrametric seminorm. To prove the reverse inequality, that is $\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}\leq\|\sum_{i\in I}x_{i}\otimes y_{i}\|_{\pi},$ we have to show that
$\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}\leq\max_{j\in J}\|x^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{K}$}},$ (60)
where $\{x^{\prime}_{j}\}$ and $\{y^{\prime}_{j}\}$ are finite sets in $\mathcal{H}$ and $\mathcal{K},$ respectively, such that $\sum_{i\in I}x_{i}\otimes y_{i}=\sum_{j\in J}x^{\prime}_{j}\otimes y^{\prime}_{j}.$ Taking the $\inf$ over all possible representations of $u$ at the r.h.s. of the inequality (60) yields the result. Let $Y\subset\mathcal{K}$ be the linear span of the $\{y_{i}\}_{i\in I}$ and fix a norm-orthogonal basis $\{e_{k}\}_{k\in K}$ for $Y$ (which it always exists, see Theorem 2.3.22 in [29]). For any element $y^{\prime}\in Y,$ we have then
$y_{i}=\sum_{k}\lambda_{ik}e_{k}\>\text{ and }\>y^{\prime}_{j}=\sum_{k}\lambda^{\prime}_{jk}e_{k}$ (61)
with $\lambda_{ik},\lambda^{\prime}_{jk}\in\mathbb{Q}_{p,\mu}.$ By linearity of the tensor product, the following chain of equations holds:
$\sum_{k\in K}\bigl(\sum_{i\in I}\lambda_{ik}x_{i}\bigr)\otimes e_{k}=\sum_{i\in I}x_{i}\otimes y_{i}=\sum_{j\in J}x^{\prime}_{j}\otimes y^{\prime}_{j}=\sum_{k\in K}\bigl(\sum_{j\in J}\lambda^{\prime}_{jk}x^{\prime}_{j}\bigr)\otimes e_{k}.$ (62)
Since the sets $\{e_{k}\}_{k\in K}$ and $\{x_{i}\}_{i\in I}$ are norm-orthogonal systems, it follows that
$\sum_{i\in I}\lambda_{ik}x_{i}=\sum_{j\in J}\lambda^{\prime}_{jk}x^{\prime}_{j},$ (63)
for all $k\in K.$ Finally we get the desired result:
$\begin{split}\max_{j\in J}\|x^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|y^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{K}$}}&\geq\max_{j\in J,k\in K}\left|\lambda^{\prime}_{jk}\right|\|x^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{H}$}}\,\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\geq\max_{k\in K}\Big(\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\cdot\|\sum_{j}\lambda^{\prime}_{jk}x^{\prime}_{j}\|_{\mbox{\tiny$\mathcal{H}$}}\Big)\\ &=\max_{k\in K}\Big(\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\,\|\sum_{i\in I}\lambda_{ik}x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\Big)\geq\max_{i\in I,k\in K}\left|\lambda_{ik}\right|\|e_{k}\|_{\mbox{\tiny$\mathcal{K}$}}\,\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\geq\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}.\end{split}$ (64)
Theorem 3.9.
The seminorm $\|\cdot\|_{\pi}$ is, actually, a norm on $\mathcal{H}\otimes_{\alpha}\mathcal{K}$, and therefore the pair $\mathcal{H}\otimes_{\pi}\mathcal{K}\equiv(\mathcal{H}\otimes_{\alpha}\mathcal{K},\|\cdot\|_{\pi})$ is a $p$-adic normed space.
Proof.
The implication $u=0\implies\|u\|_{\pi}=0$ is trivially true. For the reverse implication, we want to prove that $u\neq 0\implies\|u\|_{\pi}\neq 0$. Let $0\neq u\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$ and write it as finite sum of the kind $u=\sum_{i\in I}x_{i}\otimes y_{i}$, where $x_{i}\in\mathcal{H}$ and $y_{i}\in\mathcal{K}$ are all nonzero. Let $Y\subset\mathcal{K}$ be the linear span of the $\{y_{i}\}_{i\in I}$ and pick a norm-orthogonal basis for $Y.$ By expanding $\{y_{i}\}_{i\in I}$ with respect to this basis we can assume that it is a norm-orthogonal system in $\mathcal{K}$. Finally, Lemma 3.8 shows that $\|u\|_{\pi}=\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\cdot\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}>0.$
By taking the completion of the $p$-adic normed space $\mathcal{H}\otimes_{\pi}\mathcal{K}\equiv(\mathcal{H}\otimes_{\alpha}\mathcal{K},\|\cdot\|_{\pi})$ w.r.t. the metric induced by the norm $\|\cdot\|_{\pi}$, we obtain the $p$-adic Banach space $\big(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi}\big)$, that will be synthetically denoted by $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$.
Theorem 3.10.
If $\{x_{i}\}_{i\in I}$ is a norm-orthogonal system in $\mathcal{H}$ and $\{y_{j}\}_{j\in J}$ is a norm-orthogonal system in $\mathcal{K},$ then $\{x_{i}\otimes y_{j}\mid i\in I,\,j\in J\}$ is a norm-orthogonal system in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. In particular, if $\Phi=\{\phi_{i}\}_{i\in I}$ and $\Psi=\{\psi_{j}\}_{j\in J}$ are norm-orthogonal (normal) bases in $\mathcal{H}$ and $\mathcal{K}$, respectively, then $\bm{e}\coloneqq\{e_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}$ is a norm-orthogonal (normal) basis in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$.
Proof.
To establish the norm orthogonality of $\{x_{i}\otimes y_{j}\}_{i\in I,j\in J}$, it suffices by definition to verify it for any finite subset; hence, we may assume the index sets $I$ and $J$ to be finite. Take $\lambda_{ij}\in\mathbb{Q}_{p,\mu}$ for $i\in I$ and $j\in J$. Using Lemma 3.8 with the orthogonality of the $\{y_{j}\}_{j\in J}$ and the property that $\|x\otimes y\|_{\pi}=\|x\|_{\mbox{\tiny$\mathcal{H}$}}\cdot\|y\|_{\mbox{\tiny$\mathcal{K}$}}$ for all $x\in\mathcal{H},y\in\mathcal{K}$ we get
$\begin{split}\bigg\|\sum_{i\in I,\,j\in J}\lambda_{ij}x_{i}\otimes y_{j}\bigg\|_{\pi}&=\bigg\|\sum_{i\in I}x_{i}\otimes(\sum_{j\in J}\lambda_{ij}y_{j})\bigg\|_{\pi}=\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\,\bigg\|\sum_{j\in J}\lambda_{ij}y_{j}\bigg\|_{\mbox{\tiny$\mathcal{K}$}}\\ &=\max_{i\in I}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\max_{j\in J}\left|\lambda_{ij}\right|\,\|y_{j}\|_{\mbox{\tiny$\mathcal{K}$}}=\max_{i\in I,j\in J}\left|\lambda_{ij}\right|\|x_{i}\otimes y_{j}\|_{\pi}.\end{split}$ (65)
Now, since $\Phi$ and $\Psi$ are normal basis for $\mathcal{H}$ and $\mathcal{K}$ respectively, by Lemma 10.2.12 in [29] we have that the linear span of $\bm{e}$ is dense in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. Finally, by Theorem 9.2.1 in [29] $\bm{e}$ is therefore a norm-orthogonal basis. ∎
Eventually, we can introduce a suitable inner product on the $p$-adic Banach space $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$, in such a way as to obtain a $p$-adic Hilbert space.
Definition 3.11.
We get a non-degenerate, Hermitian sesquilinear form $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\alpha}$ on $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ extending — by linearity in the second argument and by anti-linearity in the first argument — the following definition:
$\langle x_{1}\otimes y_{1},x_{2}\otimes y_{2}\rangle_{\alpha}\coloneqq\langle x_{1},x_{2}\rangle_{\mathcal{H}}\,\langle y_{1},y_{2}\rangle_{\mathcal{K}},\quad\forall x_{1},x_{2}\in\mathcal{H},y_{1},y_{2}\in\mathcal{K}.$ (66)
The previous definition rests on the following fact:
Proposition 3.12.
There is a unique non-degenerate, Hermitian sesquilinear form $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\alpha}$ on $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ determined by (66). In fact, we have that
  1. (i)
    The quantity $\langle u,v\rangle_{\alpha}$ is well defined, that is, when computed according to Definition 3.11, it does not depend on the particular decompositions $u=\sum_{i\in I}x_{i}\otimes y_{i}$ and $v=\sum_{j\in J}x^{\prime}_{j}\otimes y^{\prime}_{j}$ of elements $u,v\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$.
  2. (ii)
    The mapping $\mathcal{H}\otimes_{\alpha}\mathcal{K}\ni(u,v)\mapsto\langle u,v\rangle_{\alpha}\in\mathbb{Q}_{p,\mu}$ is a Hermitian sesquilinear form.
  3. (iii)
    The sesquilinear form $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\alpha}$ is non-degenerate.
  4. (iv)
    The Cauchy-Schwarz inequality holds, i.e.,
    $\left|\langle u,v\rangle_{\alpha}\right|\leq\|u\|_{\mbox{\tiny$\mathcal{H}$}}\,\|v\|_{\mbox{\tiny$\mathcal{K}$}},\quad\forall u,v\in\mathcal{H}\otimes_{\alpha}\mathcal{K}.$ (67)
Proof.
To prove (i), it is sufficient to argue that, if $\sum_{j\in J}x_{j}^{\prime}\otimes y_{j}^{\prime}$ is any representation of the zero tensor $0\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, then, for any $\sum_{i\in I}x_{i}\otimes y_{i}\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, by (66) we have:
$\begin{split}\sum_{i\in I,\,j\in J}\langle x_{i}\otimes y_{i},x_{j}^{\prime}\otimes y_{j}^{\prime}\rangle_{\alpha}&=\sum_{i\in I}\sum_{j\in J}\langle x_{i},x_{j}^{\prime}\rangle_{\mathcal{H}}\,\langle y_{i},y_{j}^{\prime}\rangle_{\mathcal{K}}\\ &=0=\sum_{i\in I,\,j\in J}\langle x_{j}^{\prime}\otimes y_{j}^{\prime},x_{i}\otimes y_{i}\rangle_{\alpha}.\end{split}$ (68)
Here, the second equality follows from the implication (vii-1) $\implies$ (vii-2) in Proposition 3.6 (identifying the Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$ with their duals, in the standard fashion). At this point, by 68, one easily checks that (i) holds true.
The proof of (ii) is clear, by applying Definition 3.11.
Let us now prove (iii). In fact, given some $v=\sum_{j\in J}x_{j}^{\prime}\otimes y_{j}^{\prime}\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, if $\langle u,v\rangle_{\mathcal{K}}=0$, for every simple tensor $u=x\otimes y\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, then
$\begin{split}0&=\langle u,v\rangle_{\mathcal{K}}\\ &=\sum_{j\in J}\langle x\otimes y,x_{j}^{\prime}\otimes y_{j}^{\prime}\rangle_{\alpha}\\ &=\sum_{j\in J}\langle x,x_{j}^{\prime}\rangle_{\mathcal{H}}\,\langle y,y_{j}^{\prime}\rangle_{\mathcal{K}},\quad\forall x\in\mathcal{H},\ \forall y\in\mathcal{K};\end{split}$ (69)
hence, by the implication (vii-2) $\implies$ (vii-1) in Proposition 3.6, we conclude that $v=0$. It follows that the sesquilinear form $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\alpha}$ is non-degenerate.
Let us finally prove (iv). In fact, for every pair of vectors $u=\sum_{i\in I}x_{i}\otimes y_{i}$, $v=\sum_{j\in J}x_{j}^{\prime}\otimes y_{j}^{\prime}$ in the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}$, we have:
$\left|\langle u,v\rangle_{\alpha}\right|=\left|\sum_{i\in I}\sum_{j\in J}\langle x_{i}\otimes y_{i},x_{j}^{\prime}\otimes y_{j}^{\prime}\rangle_{\alpha}\right|$ $=\left|\sum_{i\in I}\sum_{j\in J}\langle x_{i},x_{j}^{\prime}\rangle_{\mathcal{H}}\,\langle y_{i},y_{j}^{\prime}\rangle_{\mathcal{K}}\right|$
$\leq\max_{i\in I,\,j\in J}\left|\langle x_{i},x_{j}^{\prime}\rangle_{\mathcal{H}}\right|\left|\langle y_{i},y_{j}^{\prime}\rangle_{\mathcal{K}}\right|$
$\leq\max_{i\in I,\,j\in J}\|x_{i}\|_{\mbox{\tiny$\mathcal{H}$}}\|y_{i}\|_{\mbox{\tiny$\mathcal{K}$}}\|x_{j}^{\prime}\|_{\mbox{\tiny$\mathcal{H}$}}\|y_{j}^{\prime}\|_{\mbox{\tiny$\mathcal{K}$}}=\|u\|_{\pi}\|v\|_{\pi}.$ (70)
Here, the second inequality has been obtained by the Cauchy-Schwarz inequality in the $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$.
The proof is complete. ∎
Remark 3.13.
By Proposition 3.12, $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\alpha}$ is a well-defined non-Archimedean inner product on the normed algebraic tensor product $\mathcal{H}\otimes_{\pi}\mathcal{K}\equiv(\mathcal{H}\otimes_{\alpha}\mathcal{K},\|\cdot\|_{\pi})$. In particular, it is bounded w.r.t. the ultrametric norm $\|\cdot\|_{\pi}$, by virtue of the Cauchy-Schwarz inequality (67).
Lemma 3.14.
The inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\alpha}$ on the normed algebraic tensor product $(\mathcal{H}\otimes_{\alpha}\mathcal{K},\|\cdot\|_{\pi})$ can be uniquely extended to a non-Archimedean inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$ on the $p$-adic Banach space $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. Precisely, for every pair of vectors $u,v\in\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$, we have that
$\langle u,v\rangle_{\pi}=\lim_{n\to\infty}\langle u_{n},v_{n}\rangle_{\alpha},$ (71)
where $\{u_{n}\}_{n\in\mathbb{N}}$ and $\{v_{n}\}_{n\in\mathbb{N}}$ be any two sequences in $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ such that $u_{n}\to u$ and $v_{n}\to v$ (convergence w.r.t. the norm $\|\cdot\|_{\pi}$ being understood).
Proof.
Given any pair of vectors $u,v\in\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$, let $\{u_{n}\}_{n\in\mathbb{N}}$ and $\{v_{n}\}_{n\in\mathbb{N}}$ be two sequences in $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ such that $u_{n}\to u$ and $v_{n}\to v$ (w.r.t. the norm $\|\cdot\|_{\pi}$). We first prove that the limit $\langle u,v\rangle_{\pi}\equiv\lim_{n\to\infty}\langle u_{n},v_{n}\rangle_{\alpha}$ exists and does not depend on the sequences that we have chosen. In fact, since
$|\langle u_{n},v_{n}\rangle_{\alpha}-\langle u_{m},v_{m}\rangle_{\alpha}|\leq\|u_{n}-u_{m}\|_{\pi}\|v_{n}\|_{\pi}+\|u_{m}\|_{\pi}\|v_{n}-v_{m}\|_{\pi},$ (72)
we have that $\{\langle u_{n},v_{n}\rangle_{\alpha}\}_{n\in\mathbb{N}}$ is a Cauchy sequence and, therefore, is convergent. To prove that its limit is unique, let us take any other pair of sequences $\{u_{n}^{\prime}\}_{n\in\mathbb{N}}$ and $\{v_{n}^{\prime}\}_{n\in\mathbb{N}}$ such that $u_{n}^{\prime}\to u$ and $v_{n}^{\prime}\to v$. We have:
$|\langle u_{n},v_{n}\rangle_{\alpha}-\langle u_{n}^{\prime},v_{n}^{\prime}\rangle_{\alpha}|\leq\|u_{n}-u_{n}^{\prime}\|_{\pi}\|v_{n}\|_{\pi}+\|u_{n}^{\prime}\|_{\pi}\|v_{n}-v_{n}^{\prime}\|_{\pi}\to 0.$ (73)
By the previous two points, for every pair of vectors $u,v\in\mathcal{H}\otimes_{\alpha}\mathcal{K}$, considering the sequences $\{u,u,\ldots\}$ and $\{v,v,\ldots\}$, we see that $\langle u,v\rangle_{\alpha}=\langle u,v\rangle_{\pi}$.
Moreover, one can easily check that the mapping $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}\ni u,v\mapsto\langle u,v\rangle_{\pi}\in\mathbb{Q}_{p,\mu}$ is a Hermitian sesquilinear form, that, by the previous point, extends $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\alpha}$ to the Banach space completion $(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi})$ of the normed algebraic tensor product $(\mathcal{H}\otimes_{\alpha}\mathcal{K},\|\cdot\|_{\pi})$.
Finally, let us observe that
$\left|\langle u,v\rangle_{\pi}\right|=\lim_{n\to\infty}\left|\langle u_{n},v_{n}\rangle_{\alpha}\right|\leq\lim_{n\to\infty}\|u_{n}\|_{\pi}\|v_{n}\|_{\pi}=\|u\|_{\pi}\|v\|_{\pi},$ (74)
which proves that the Hermitian sesquilinear form $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$ satisfies the Cauchy-Schwarz inequality too, hence, it is a non-Archimedean inner product on the $p$-adic Banach space $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. Clearly, every sesquilinear form, that extends $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\alpha}$ to the Banach space completion of the normed algebraic tensor product $(\mathcal{H}\otimes_{\alpha}\mathcal{K},\|\cdot\|_{\pi})$ and satisfies the Cauchy-Schwarz inequality, is continuous and, therefore, is uniquely determined by a condition of the form (71); otherwise stated, it must coincide with $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$. ∎
Theorem 3.15.
The quadruple $\big(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi},\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi},\bm{e}\equiv\{e_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}\big)$ — which, in the following, will be synthetically denoted by $\mathcal{H}\otimes\mathcal{K}$ — is a $p$-adic Hilbert space.
Proof.
The pair $(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi})$ is a $p$-adic Banach space that, further endowed with the non-Archimedean inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$ defined by relation (71) (Lemma 3.14), becomes a inner product $p$-adic Banach space. Finally, to conclude our construction, we observe that the norm-orthogonal basis $\bm{e}\equiv\{e_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}$ in $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$ is also IP-orthogonal w.r.t. the inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$, and, moreover, $\langle e_{k},e_{l}\rangle_{\pi}=\delta_{kl}$. ∎
Remark 3.16.
It should be clear that the synthetic symbol $\mathcal{H}\otimes\mathcal{K}$ denoting the tensor product of the $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$ (i.e., a $p$-adic Hilbert space itself) mimics the usual symbol $\mathscr{H}\otimes\mathscr{K}$ for the tensor product of two complex Hilbert spaces $\mathscr{H}$ and $\mathscr{K}$.
Corollary 3.17.
Given orthonormal bases $\{\phi_{i}\}_{i\in I}$ and $\{\psi_{j}\}_{j\in J}$ in $\mathcal{H}$ and $\mathcal{K}$, respectively, let $\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}$ be the associated orthonormal basis in the $p$-adic Hilbert space $\mathcal{H}\otimes\mathcal{K}$. Then, for every vector $u=\sum_{i\in I,\,j\in J}\lambda_{ij}\,\phi_{i}\otimes\psi_{j}\in\mathcal{H}\otimes\mathcal{K}$, we have that $\lambda_{ij}=\langle\phi_{i}\otimes\psi_{j},u\rangle_{\pi}$ and
$\|u\|_{\pi}=\max_{i\in I,\,j\in J}|\lambda_{ij}|.$ (75)
Proof.
This result is directly derived from Theorem 3.15 and the non-Archimedean Parseval identity in [22]. ∎
4 Relation with $p$-adic Hilbert-Schmidt-class operators
As is well known, there exists an Hilbert space isomorphism between the tensor product of (complex) Hilbert spaces $\mathscr{H}\otimes\mathscr{K}$, and the space of Hilbert-Schmidt operators from $\mathscr{H}$ to $\mathscr{K}$ [36, 37, 38]. Here we show that this picture admits a natural extension to the $p$-adic setting; namely, we prove that there exists an inner-product preserving surjective isometry — i.e. an isomorphism of $p$-adic Hilbert spaces — between $\mathcal{T}(\mathcal{H},\mathcal{K})$ and $\mathcal{H}\otimes\mathcal{K}$. However, as a preliminary step, we need to first introduce a suitable notion of antiunitary operator in a $p$-adic Hilbert space.
4.1 Anti-unitary operators
We start with the following
Definition 4.1.
We say that an operator $A$ in a $p$-adic Hilbert space $\mathcal{H}$ is conjugate-linear (or, anti-linear), if the relation
$A(\alpha\psi+\beta\phi)=\overline{\alpha}A\phi+\overline{\beta}A\psi$ (76)
holds for every $\alpha,\beta\in\mathbb{Q}_{p,\mu}$, and $\phi,\psi\in\mathrm{dom}(A)$.
Similar to linear operators, a conjugate-linear operator $A\colon\mathcal{H}\dashrightarrow\mathcal{H}$ is continuous precisely when it is bounded; moreover, we say that $A$ is all-over if $\mathrm{dom}(A)=\mathcal{H}$.
Definition 4.2.
Given any orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$, a bounded conjugate-linear operator $J_{\Phi}\colon\mathcal{H}\rightarrow\mathcal{H}$ such that
$J_{\Phi}\,\phi_{m}=\phi_{m},\quad\forall m\in\mathbb{N},$ (77)
is called a conjugation operator. In this case, we will say that $J_{\Phi}$ is the conjugation operator associated with the orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$.
Remark 4.3.
Given any orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$, it is always possible to construct a conjugation operator $J_{\Phi}$ associated with $\Phi$ (see the discussion following Example 4.9). Moreover, it is not difficult to verify that if $L$ is a linear operator in $\mathcal{H}$, the operator defined as $A\coloneqq J_{\Phi}L$ (or, equivalently, $A\coloneqq LJ_{\Phi}$), is a conjugate-linear operator and, conversely, any conjugate-linear operator $A$ in $\mathcal{H}$ can be expressed in this form, for some linear operator $L$ in $\mathcal{H}$.
Remark 4.4.
We recall that a linear operator $L\colon\mathcal{H}\dashrightarrow\mathcal{H}$, in a $p$-adic Hilbert space $\mathcal{H}$, is norm-preserving (N-preserving, in short), if it satisfies the condition: $\|L\phi\|=\|\phi\|$, for any $\phi\in\mathrm{dom}(L)$ [22]. An all-over N-preserving map in $\mathcal{H}$ is a linear isometry of $\mathcal{H}$. A linear operator $L$ in $\mathcal{H}$ is inner-product-preserving (IP-preserving), if
$\langle L\phi,L\psi\rangle=\langle\phi,\psi\rangle,\quad\forall\phi,\psi\in\mathrm{dom}(L),$ (78)
while $L$ is norm-orthogonality-preserving (NO-preserving), if it is linear, and satisfies the condition [22]:
$\phi,\psi\in\mathrm{dom}(L),\quad\phi\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}\psi\;\implies\;L\phi\overset{\mbox{\tiny$\|\hskip-1.42262pt\cdot\hskip-1.42262pt\|$}}{\perp}L\psi.$ (79)
An IP-preserving surjective isometry (i.e., an automorphism) of $\mathcal{H}$ is a unitary operator; adopting the notation already used in [22], in the following we will denote the set of all such operators by $\mathcal{U}(\mathcal{H})$. (For a complete characterization of unitary operators in a $p$-adic Hilbert space see, in particular, Theorem $5.13$ in [22]).
Bearing in mind Remarks 4.3 and 4.4, it is now natural to set the following
Definition 4.5.
Let $A\colon\mathcal{H}\dashrightarrow\mathcal{H}$ be a conjugate-linear operator in a $p$-adic Hilbert space $\mathcal{H}$. We say that
  1. (A1)
    AA is an anti-isometric operator, if $A$ is all-over and N-preserving — i.e., if $\|A\phi\|=\|\phi\|$ for every $\phi\in\mathrm{dom}(A)=\mathcal{H}$.
  2. (A2)
    AA is a conjugate-linear topological isomorphism, if $A$ is bounded and admits a (conjugate-linear) bounded inverse.
  3. (A3)
    AA is inner-product-conjugating, if the condition
    $\langle A\phi,A\psi\rangle=\overline{\langle\phi,\psi\rangle}$ (80)
    is satisfied for every $\phi,\psi$ in $\mathrm{dom}(A)$.
Proposition 4.6.
Let $\mathcal{H}$ be a $p$-adic Hilbert space, let $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ be an orthonormal basis, and let $J_{\Phi}$ be the associated conjugation operator. Then, the following facts hold:
  1. (i)
    For every vector $\chi$ in $\mathcal{H}$, we have:
    $J_{\Phi}\chi=J_{\Phi}\bigg(\sum_{n}\langle\phi_{n},\chi\rangle\phi_{n}\bigg)=\sum_{n}\overline{\langle\phi_{n},\chi\rangle}\phi_{n}.$ (81)
    Moreover, $J_{\Phi}$ is IP-conjugating and anti-isometric.
  2. (ii)
    $J_{\Phi}$ is involutive, i.e., $J_{\Phi}^{2}=\mathrm{Id}$, bijective, bounded, and adjointable. In particular, $J_{\Phi}=J_{\Phi}^{-1}=J_{\Phi}^{\ast}$.
Proof.
Relation (81) simply follows by anti-linearity from Definition 4.2. Then, we have:
$\langle J_{\Phi}\chi,J_{\Phi}\eta\rangle=\Big\langle\sum_{n}\overline{\langle\phi_{n},\chi\rangle}\phi_{n},\,\sum_{m}\overline{\langle\phi_{m},\eta\rangle}\phi_{m}\Big\rangle=\sum_{n}\langle\phi_{n},\chi\rangle\overline{\langle\phi_{m},\eta\rangle}=\overline{\langle\chi,\eta\rangle},$ (82)
i.e., $J_{\Phi}$ is IP-conjugating. Moreover, by observing that
$\|J_{\Phi}\chi\|=\|\mbox{$\sum_{n}\overline{\langle\phi_{n},\chi\rangle}\phi_{n}$}\|=\sup_{n}|\overline{\langle\phi_{n},\chi\rangle}|=\sup_{n}|\langle\phi_{n},\chi\rangle|=\|\chi\|,$ (83)
and since, from (81), $\mathrm{dom}(J_{\Phi})=\mathcal{H}$, we conclude that $J_{\Phi}$ is an anti-isometric operator, i.e., (i) holds. Let us now prove (ii). Exploiting (81) twice, we have:
$J_{\Phi}^{2}\chi=J_{\Phi}(J_{\Phi}\chi)=J_{\Phi}\bigg(\sum_{n}\overline{\langle\phi_{n},\chi\rangle}\phi_{n}\bigg)=\sum_{n}\langle\phi_{n},\chi\rangle\phi_{n},$ (84)
i.e., $J_{\Phi}^{2}=\mathrm{Id}$. Now, it is clear that $J_{\Phi}$ is surjective and, by point (i), an anti-isometric operator. Therefore, $J_{\Phi}$ is bijective, bounded, and $J_{\Phi}^{-1}=J_{\Phi}$. Moreover, by observing that
$\langle\phi_{n},J_{\Phi}^{\ast}\phi_{m}\rangle=\overline{\langle J_{\Phi}^{-1}\phi_{n},J_{\Phi}^{\ast}J^{-1}_{\Phi}\phi_{m}\rangle}=\overline{\langle J_{\Phi}J_{\Phi}^{-1}\phi_{n},J_{\Phi}^{-1}\phi_{m}\rangle}=\overline{\langle\phi_{n},\phi_{m}\rangle}=\delta_{nm},$ (85)
we also see that $J_{\Phi}^{\ast}\phi_{m}=\phi_{m}$, for all $m\in\mathbb{N}$, i.e., $J_{\Phi}^{\ast}=J_{\Phi}$. ∎
As pointed out in Remark 4.3, composing a linear operator $L$ in $\mathcal{H}$ with the conjugation operator $J_{\Phi}$ — associated with an orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}\subset\mathcal{H}$ — yields a conjugate-linear operator in $\mathcal{H}$; conversely, every conjugate-linear operator in $\mathcal{H}$ can be expressed in this form. With this in mind, we are naturally led to the following
Definition 4.7.
Let $\mathcal{H}$ be a $p$-adic Hilbert space, and let $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ be an orthonormal basis. We say that $Z\colon\mathcal{H}\rightarrow\mathcal{H}$ is an anti-unitary operator, if it can be expressed in the form
$Z=J_{\Phi}U,\quad U\in\mathcal{U}(\mathcal{H}),$ (86)
i.e., as the composition of the conjugation operator $J_{\Phi}$ — associated with the orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$ — and some unitary operator $U\in\mathcal{U}(\mathcal{H})$. We will denote the set of all the anti-unitary operators in $\mathcal{H}$ by $\overline{\mathcal{U}}(\mathcal{H})$.
The next result provides a complete characterization of anti-unitary operators.
Theorem 4.8.
Given a conjugate-linear operator $Z$ in $\mathcal{H}$, the following facts are equivalent:
  1. (Z1)
    $Z$ is an anti-unitary operator, i.e., $Z=J_{\Phi}U$, for $J_{\Phi}$ the conjugation operator associated with an orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$, and $U$ some unitary operator in $\mathcal{U}(\mathcal{H})$;
  2. (Z2)
    $Z$ is bounded and, for some pair of orthonormal bases $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$, we have that $Z\phi_{k}=\psi_{k}$;
  3. (Z3)
    $Z$ is bounded and adjointable, $ZZ^{\ast}=Z^{\ast}Z=\mathrm{Id}$, and $\|Z\|=1$;
  4. (Z4)
    $Z$ is a surjective, IP-conjugating, all-over operator, and $\|Z\|=1$;
  5. (Z5)
    $Z$ is a IP-conjugating, conjugate-linear topological isomorphism, and $\|Z\|=1=\|Z^{-1}\|$;
  6. (Z6)
    $Z$ is an anti-automorphism of the $p$-adic Hilbert space $\mathcal{H}$, namely, a IP-conjugating surjective anti-isometry;
  7. (Z7)
    $Z$ is a surjective, IP-conjugating, NO-preserving, all-over operator;
  8. (Z8)
    $Z$ is bounded and transforms orthonormal bases into orthonormal bases.
Proof.
Let us first prove (Z1) $\iff$ (Z2) $\implies$ (Z3). Indeed, let $Z\in\overline{\mathcal{U}}(\mathcal{H})$ and let $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ be two orthonormal bases in $\mathcal{H}$. Let $J_{\Psi}$ be the conjugation operator associated with $\Psi$ — i.e., $J_{\Psi}\psi_{n}=\psi_{n}$, $\forall n\in\mathbb{N}$. Since $Z$ is anti-unitary, there exists a unitary operator $U$ in $\mathcal{U}(\mathcal{H})$ such that $Z=J_{\Psi}U$. Thus, we have:
$Z\phi_{k}=J_{\Psi}U\phi_{k}=J_{\Psi}\psi_{k}=\psi_{k},\quad\forall k\in\mathbb{N},$ (87)
where we have used the fact that since $U$ is unitary, it maps orthonormal bases into orthonormal bases (cf. point (U22) of Theorem $5.13$ in [22]). Moreover, it is clear that $Z=J_{\Psi}U$ is bounded; hence, (Z1) $\implies$ (Z2). Conversely, let $Z$ be as in (Z2). Since it is conjugate-linear, we can express $Z$ as $Z=J_{\Psi}L$, where $J_{\Psi}$ is the conjugation operator associated with the orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$, and $L$ is some linear operator in $\mathcal{H}$ (see Remark 4.3). Now, since, by assumption, $Z$ is a bounded operator, $L$ is bounded too. Moreover, by observing that
$\psi_{k}=Z\phi_{k}=J_{\Psi}L\phi_{k}=\psi_{k}=J_{\Psi}\psi_{k},\quad\forall k\in\mathbb{N},$ (88)
we can conclude that $L\phi_{k}=\psi_{k}$, for all $k\in\mathbb{N}$. Hence, exploiting point (U22) of Theorem $5.13$ in [22], $L$ is unitary and, therefore, $Z$ is an anti-unitary operator. We next prove (Z2) $\implies$ (Z3). We have already shown that an anti-unitary operator is completely characterized by condition (Z2). Now, we have that
$\langle\phi_{m},Z^{\ast}\psi_{k}\rangle=\langle Z\phi_{m},\psi_{k}\rangle=\langle J_{\Psi}U\phi_{m},\psi_{k}\rangle=\langle J_{\Psi}\psi_{m},\psi_{k}\rangle=\delta_{mk},$ (89)
where we have expressed $Z$ as $Z=J_{\Psi}U$, and used the fact that, since $U$ is unitary, it maps the orthonormal basis $\Phi=\{\phi_{m}\}_{m\in\mathbb{N}}$ into $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$. Therefore, $Z$ is adjointable and $Z^{\ast}\psi_{k}=\phi_{k}$, $\forall k\in\mathbb{N}$, i.e., $Z^{\ast}$ is anti-unitary. From this, it also follows that
$Z^{\ast}Z=\mathrm{Id}=ZZ^{\ast}.$ (90)
Since $Z=J_{\Psi}U$ is the composition of the anti-isometric operator $J_{\Psi}$, and the unitary operator $U$, we observe that
$\|Z\|=\|J_{\Psi}U\|=\sup_{\chi\neq 0}\frac{\|J_{\Psi}U\chi\|}{\|\chi\|}=\sup_{\chi\neq 0}\frac{\|U\chi\|}{\|\chi\|}=1,$ (91)
and, hence,
$\|Z\|=\|Z^{\ast}\|=1.$ (92)
The implication (Z3) $\implies$ (Z4) is clear since, if $Z$ bounded, adjointable, and satisfies condition (90), then $Z$ is surjective and all-over. Moreover, if $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ is an orthonormal basis in $\mathcal{H}$, by observing that
$\langle Z\chi,Z\eta\rangle$ $=\Big\langle Z\sum_{n}\langle\psi_{n},\chi\rangle\psi_{n},\sum_{m}\langle\psi_{m},\eta\rangle\psi_{m}\Big\rangle$
$=\Big\langle\sum_{n}\overline{\langle\chi,\psi_{n}\rangle}Z\psi_{n},\sum_{m}\overline{\langle\psi_{m},\eta\rangle}Z\psi_{m}\Big\rangle$
$=\sum_{n}\sum_{m}\langle\psi_{n},\chi\rangle\overline{\langle\psi_{m},\eta\rangle}\langle Z\psi_{n},Z\psi_{m}\rangle$
$=\sum_{n}\sum_{m}\langle\psi_{n},\chi\rangle\overline{\langle\psi_{m},\eta\rangle}\langle Z^{\ast}Z\psi_{n},\psi_{m}\rangle$
$=\sum_{n}\langle\psi_{n},\chi\rangle\overline{\langle\psi_{n},\eta\rangle}=\overline{\langle\chi,\eta\rangle},$ (93)
for all $\chi,\eta$ in $\mathcal{H}$, we see that $Z$ is IP-conjugating too and, hence, (Z4) is verified. Next, let us prove (Z4) $\implies$ (Z5). Expressing $Z$ as $Z=J_{\Psi}L$ — for a suitable linear operator $L$ in $\mathcal{H}$, and with $J_{\Psi}$ the conjugation operator associated with the orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ — since, by assumption, $Z$ is surjective and IP-conjugating, we see that $L$ must be surjective and IP-preserving. Then, Theorem $5.8$ in [22] entails that $L$ is an adjointable top-linear isomorphism such that $L^{\ast}=L^{-1}$ and $\|L\|=\|L^{\ast}\|=\|L^{-1}\|$. This then yields
$L^{\ast}J_{\Psi}=L^{-1}J_{\Psi}\;\;\implies\;\;(J_{\Psi}L)^{\ast}=(J_{\Psi}L)^{-1},$ (94)
i.e., $Z^{\ast}=Z^{-1}$. (Note that, in the second equality in (94), we have used property (ii) of Proposition 4.6). Hence, $Z$ is an adjointable conjugate-linear topological isomorphism and
$\|Z\|=\|Z^{\ast}\|=\|Z^{-1}\|=1.$ (95)
Now, if $Z$ satisfies (Z5), expressing again $Z$ as $Z=J_{\Psi}L$, for some linear operator $L$, we see that, by hypothesis, $L$ must be an IP-preserving top-linear isomorphism such that $\|L\|=\|L^{-1}\|$. Then, by Lemma $5.2$ in [22], $L$ is an IP-reserving surjective isometry. Therefore, $Z=J_{\Psi}L$ is an IP-conjugating surjective anti-isometry, i.e., (Z5) $\implies$ (Z6). Let us now show (Z6) $\iff$ (Z7). Indeed, if $Z=J_{\Psi}L$ is anti-isometric, then $L$ is an isometry, and, hence, by Theorem $5.11$ in [22], an NO-preserving all-over operator. Taking into account that $J_{\Psi}$ is an anti-isometry, we see that $Z$ is an IP-conjugating, surjective, NO-preserving, all-over operator, i.e., (Z7) holds. Conversely, assume that $Z=J_{\Psi}L$ is IP-conjugating, surjective and all-over. Then, $L$ is a surjective IP-preserving operator and, hence, $\|L\|=\|L^{-1}\|$ (see Theorem $5.8$ in [22]). From this, it follows that
$\|Z\|=\|J_{\Psi}L\|=\|L\|=\|L^{-1}\|=\|L^{-1}J_{\Psi}^{-1}\|=\|Z^{-1}\|.$ (96)
Next, let us further assume that $Z$ is NO-preserving. Since $J_{\Psi}$ is an anti-isometry, it must be true that $L$ is an all-over and NO-preserving linear operator; then, by Theorem $5.11$ in [22], $L$ must be a non-zero scalar multiple of an isometry, i.e., $L=zI$, where $z\in\mathbb{Q}_{p,\mu}\setminus\{0\}$, and $I\colon\mathcal{H}\rightarrow\mathcal{H}$ is an isometry in $\mathcal{H}$. We thus have that $Z=J_{\Psi}zI=\overline{z}J_{\Psi}I$. From this, and owing to (96), we see that
$\left|z\right|=\left|\overline{z}\right|=\|J_{\Psi}Iz\|=\|Z\|=\|Z^{-1}\|=\|\overline{z}^{-1}I^{-1}J_{\Psi}^{-1}\|=\left|\overline{z}\right|^{-1}=\left|z\right|^{-1},$ (97)
i.e., $\left|z\right|=1$ and $Z=\overline{z}J_{\Psi}I$ is a surjective IP-conjugating anti-isometry, i.e., (Z6) holds. To complete the proof, it is left to show that (Z6) $\implies$ (Z8) and (Z8) $\implies$ (Z2). That (Z8) $\implies$ (Z2) is clear. Hence, let us prove (Z6) $\implies$ (Z8). Assume that $Z$ satisfies conditions in (Z6), i.e., $Z$ is an IP-conjugating surjective anti-isometry. Now, given any orthonormal basis $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$, it is readily seen that the set $\{Z\phi_{n}\}_{n\in\mathbb{N}}$ provides another orthonormal basis in $\mathcal{H}$. Indeed, since $W$ is IP-conjugating, we have:
$\langle Z\phi_{m},Z\phi_{n}\rangle=\overline{\langle\phi_{n},\phi_{m}\rangle}=\delta_{mn},$ (98)
i.e., $\{Z\phi_{n}\}_{n\in\mathbb{N}}$ is an IP-orthonormal set; moreover, for every finite or denumerable set $\{\alpha_{n}\}_{n\in\mathbb{N}}\subset\mathbb{Q}_{p,\mu}$, and every $\eta$ in $\mathcal{H}$, we have:
$\|\mbox{$\sum_{n}z_{n}Z\phi_{m}$}\|=\|\mbox{$Z\sum_{n}\overline{z}_{n}\phi_{m}$}\|=\max_{n}\left|\overline{z}_{n}\right|=\max_{n}\left|z_{n}\right|,$ (99)
namely, $\{Z\phi_{n}\}_{n\in\mathbb{N}}$ is a normal set. Eventually by observing that
$\chi=Z(Z^{-1}\chi)$ $=Z\sum_{n}\langle\phi_{n},Z^{-1}\chi\rangle\phi_{n}=\sum_{n}\overline{\langle\phi_{n},Z^{-1}\chi\rangle}Z\phi_{n}=$
$=\sum_{n}\langle Z\phi_{n},Z^{-1}Z\chi\rangle Z\phi_{n}=\sum_{n}\langle Z\phi_{n},\chi\rangle Z\phi_{n},$ (100)
we see that $\{Z\phi_{n}\}_{n\in\mathbb{N}}$ is dense in $\mathcal{H}$ and, hence, an orthonormal basis. This then proves that (Z6) $\implies$ (Z8) and, therefore, concludes the proof. ∎
As is well known, in a complex Hilbert space $\mathscr{K}$, every involutive, anti-unitary operator acts as a conjugation operator w.r.t. some orthonormal basis in $\mathscr{K}$ [28]. Specifically, if $Z$ is anti-unitary and satisfies $Z^{2}=\mathrm{Id}$, then an orthonormal basis $\Xi\equiv\{\xi_{k}\}_{k\in\mathbb{N}}$ in $\mathscr{K}$ exists such that $Z\xi_{k}=\xi_{k}$, for all $k\in\mathbb{N}$ — i.e., $Z\equiv J_{\Xi}$ is the conjugation operator associated with $\Xi\equiv\{\xi_{k}\}_{k\in\mathbb{N}}$ (see also Remark 4.11). In the next example, we show that, given an involutive anti-unitary operator $Z$ in a $p$-adic Hilbert space $\mathcal{H}$, one cannot always come up with an orthonormal basis $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$ preserved by $Z$.
Example 4.9.
Let $\mathcal{H}$ be a $p$-adic Hilbert space, and let $\Phi\equiv\{\phi_{m}\}_{m\in\mathbb{N}}$ be an orthonormal basis. Let $Z\in\overline{\mathcal{U}}(\mathcal{H})$ be such that $Z^{2}=\mathrm{Id}$ (i.e., $Z$ is an involutive anti-unitary operator in $\mathcal{H}$). Define the set of vectors $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ in $\mathcal{H}$ by setting
$\psi_{2m}=z_{1}\big(Z\phi_{m}+\phi_{m}\big),\quad\psi_{2m-1}=\frac{z_{2}}{\sqrt{\mu}}\big(Z\phi_{m}-\phi_{m}\big),$ (101)
where $z_{1},z_{2}$ are suitably chosen coefficients in $\mathbb{Q}_{p}$. Note that, since $Z$ is anti-unitary, $\{Z\phi_{m}\}_{m\in\mathbb{N}}$ is an orthonormal set in $\mathcal{H}$ (see (Z8) in Theorem 4.8). By construction, $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is a $Z$-invariant dense subset of $\mathcal{H}$ — just note that $\phi_{m}=\frac{1}{2}z_{1}\psi_{2m}-\frac{\sqrt{\mu}}{2}z_{2}\psi_{2m-1}$. By suitably choosing the coefficients $z_{1},z_{2}$ in $\mathbb{Q}_{p}$, we can actually turn $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ into a normal set. Specifically, let $z_{1}$ and $z_{2}$ be fixed in such a way that
$\max\big\{\left|1-z_{1}\right|,\left|z_{1}\right|\big\}\equiv t_{1}<1,\;\;\mbox{and}\;\;\max\bigg\{\left|1+\frac{z_{2}}{\sqrt{\mu}}\right|,\left|\frac{z_{2}}{\sqrt{\mu}}\right|\bigg\}\equiv t_{2}<1.$ (102)
With this choice of the coefficients $z_{1}$ and $z_{2}$, it is clear that $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is a normal basis in $\mathcal{H}$. Indeed, it suffices to notice that, for $m\in\mathbb{N}$ even, we have:
$\|\phi_{m}-\psi_{m}\|=\|\phi_{m}(1-z_{1})-z_{1}Z\phi_{m}\|\leq\max\big\{\left|1-z_{1}\right|,\left|z_{1}\right|\big\}\|\phi_{m}\|=t_{1}\|\phi_{m}\|<\|\phi_{m}\|=1,$ (103)
where, in the first inequality, we have used the fact that since $Z$ is an anti-unitary operator, it is isometric, i.e., $\|Z\phi_{m}\|=\|\phi_{m}\|$. Similarly, for $m\in\mathbb{N}$ odd, we find
$\|\phi_{m}-\psi_{m}\|=\bigg\|\bigg(1+\frac{z_{2}}{\sqrt{\mu}}\bigg)\phi_{m}-\frac{z_{2}}{\sqrt{\mu}}Z\phi_{m}\bigg\|\leq\max\bigg\{\left|1+\frac{z_{2}}{\sqrt{\mu}}\right|,\left|\frac{z_{2}}{\sqrt{\mu}}\right|\bigg\}\|\phi_{m}\|=t_{2}\|\phi_{m}\|<1.$ (104)
Therefore, setting $t\equiv\max\{t_{1},t_{2}\}$, from (103) and (104) and using the well known Full Perturbation Theorem for $t$-orthogonal bases in a $p$-adic normed space (see Theorem $2.3.16$ in [29]), we can argue that $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is a normal basis in $\mathcal{H}$. Next, let us see the constraints that $z_{1},z_{2}$ have to satisfy for $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ to be IP-orthonormal too. It is clear that $\langle\psi_{n},\psi_{m}\rangle=0$, for all $n\neq m\in\mathbb{N}$, i.e., $\Psi$ is IP-orthogonal; requiring further the IP-orthonormality we get the following conditions:
$1=\langle\psi_{m},\psi_{m}\rangle=\langle z_{1}(Z\phi_{m}+\phi_{m}),z_{1}(Z\phi_{m}+\phi_{m})\rangle=2z_{1}^{2}$ (105)
for $m$ even, while, for $m$ odd
$1=\langle\psi_{m},\psi_{m}\rangle=\bigg\langle\frac{z_{2}}{\sqrt{\mu}}\big(Z\phi_{m}-\phi_{m}\big),\frac{z_{2}}{\sqrt{\mu}}\big(Z\phi_{m}-\phi_{m}\big)\bigg\rangle=-\frac{2z_{2}^{2}}{\mu}.$ (106)
Therefore, (105) and (106) entail that
$z_{1}=\sqrt{\frac{1}{2}},\quad\mbox{and},\quad z_{2}=\sqrt{-\frac{\mu}{2}},$ (107)
i.e., $z_{1}$, $z_{2}$ do not belong to $\mathbb{Q}_{p}$. On the other hand, if, in the defining condition of the set $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ in (101) above, we assume that $z_{1},z_{2}$ are in $\mathbb{Q}_{p,\mu}$ — i.e., $z_{1}$ and $z_{2}$ belong to a suitable quadratic extension of $\mathbb{Q}_{p}$ — setting $z_{1}$ and $z_{2}$ as in (107), we can turn $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ into an orthonormal basis; but, in this case, it will not be an invariant set under the action of $Z$. Concluding, we have the following dichotomy: Either $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is a $Z$-invariant, normal, IP-orthogonal basis — in which case it is not IP-orthonormal — or $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ is an orthonormal basis, but not $Z$-invariant.
In example 4.9 we construct an invariant dense subset of $\mathcal{H}$, which cannot be turned into an IP-orthonormal set. Actually, this example is a manifestation of a peculiar feature of the conjugation operators in the $p$-adic setting, as we will see in the next Proposition 4.10. Before stating this result, let us first note the following. Consider the canonical Hilbert space $\mathbb{H}$ endowed with the (so-called) canonical basis $\bm{e}\equiv\{e_{i}\}_{i\in\mathbb{N}}$ (see Example 2.11). Define in $\mathbb{H}$ the transformation $J_{0}$ by setting
$J_{0}\coloneqq\mathcal{P}^{-1}\overline{(\,\cdot\,)}\mathcal{P},$ (108)
where $\mathcal{P}\colon\mathbb{H}\rightarrow c_{0}(\mathbb{N},\mathbb{Q}_{p,\mu})$ is the map defined as
$\mathbb{H}\ni\xi=\mbox{$\sum_{i}\xi_{i}e_{i}$}\mapsto\mathcal{P}(\xi)\coloneqq\{\xi_{i}\}_{i\in\mathbb{N}}\in c_{0}(\mathbb{N},\mathbb{Q}_{p,\mu}),$ (109)
while $\overline{(\,\cdot\,)}\colon c_{0}(\mathbb{N},\mathbb{Q}_{p,\mu})\ni\{\xi_{i}\}_{i\in\mathbb{N}}\mapsto\{\overline{\xi}_{i}\}_{i\in\mathbb{N}}\in c_{0}(\mathbb{N},\mathbb{Q}_{p,\mu})$ is the conjugate-linear isometry of $c_{0}(\mathbb{N},\mathbb{Q}_{p,\mu})$ into itself. It is clear that $J_{0}$ is anti-unitary and involutive. Moreover, by observing that, for every $\xi=\sum_{i}\xi_{i}e_{i}$ in $\mathbb{H}$,
$J_{0}\xi=\mathcal{P}^{-1}\overline{(\,\cdot\,)}\mathcal{P}\Big(\mbox{$\sum_{i}\xi_{i}e_{i}$}\Big)=\mathcal{P}^{-1}(\{\overline{\xi}_{i}\}_{i\in\mathbb{N}})=\mbox{$\sum_{i}\overline{\xi}_{i}e_{i}$},$ (110)
we can argue that $J_{0}$ acts as the conjugation operator associated with the canonical basis $\{e_{i}\}_{i\in\mathbb{N}}$ of $\mathbb{H}$ (in the following, we will refer to $J_{0}$ as to the canonical conjugation in $\mathbb{H}$). Let now $\mathcal{H}$ be a $p$-adic Hilbert space, and let $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ be an orthonormal basis. Let $W_{\Psi}$ be the canonical isomorphism of $\mathcal{H}$ into $\mathbb{H}$ (see (14)). It is clear that the (conjugate-linear) operator defined as
$J_{\Psi}\coloneqq W_{\Psi}^{-1}J_{0}W_{\Psi},$ (111)
provides the conjugation operator associated with the orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$. Indeed, $J_{\Psi}$ is anti-unitary and involutive; moreover, for every $\chi=\sum_{n}\langle\psi_{n},\chi\rangle\psi_{n}$ in $\mathcal{H}$, we have:
$J_{\Psi}\chi$ $=J_{\Psi}\bigg(\sum_{n}\langle\psi_{n},\chi\rangle\psi_{n}\bigg)=W_{\Psi}^{-1}J_{0}W_{\Psi}\bigg(\sum_{n}\langle\psi_{n},\chi\rangle\psi_{n}\bigg)$
$=W_{\Psi}^{-1}J_{0}\bigg(\sum_{n}\langle\psi_{n},\chi\rangle e_{n}\bigg)=W_{\Psi}^{-1}\bigg(\sum_{n}\overline{\langle\psi_{n},\chi\rangle}J_{0}e_{n}\bigg)$
$=W_{\Psi}^{-1}\bigg(\sum_{n}\overline{\langle\psi_{n},\chi\rangle}e_{n}\bigg)=\sum_{n}\overline{\langle\psi_{n},\chi\rangle}\psi_{n},$ (112)
i.e., $J_{\Psi}\psi_{n}=W_{\Psi}^{-1}J_{0}W_{\Psi}\psi_{n}=\psi_{n}$. We have thus proved that, given any orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$, it is always possible to construct a conjugation operator associated with it; moreover, the so-constructed conjugation operator is equivalent, up to isomorphisms, with the conjugation $J_{0}$ in the canonical Hilbert space $\mathbb{H}$. A natural question to ask is whether the reverse implication holds, that is, if, given an involutive anti-unitary operator $Z$ in $\mathcal{H}$, there always exists an Hilbert space isomorphism mapping $Z$ into the canonical conjugation operator $J_{0}$. The next result shows that, in fact, it is not so.
Proposition 4.10.
Let $Z$ be an involutive anti-unitary operator in the $p$-adic Hilbert space $\mathcal{H}$ — i.e., $Z\in\overline{\mathcal{U}}(\mathcal{H})$, and $Z^{2}=\mathrm{Id}$. Then, in general, there exists no isomorphism $\mathscr{I}\colon\mathbb{H}\rightarrow\mathcal{H}$, such that $\mathscr{I}^{-1}Z\mathscr{I}=J_{0}$, where $J_{0}$ is the conjugation operator associated with the canonical basis $\bm{e}\equiv\{e_{i}\}_{i\in\mathbb{N}}$ in the coordinate Hilbert space $\mathbb{H}$.
Proof.
The existence of an isomorphism as stated in the proposition is equivalent to requiring the existence of a $Z$-invariant orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$. Indeed, if an isomorphism $\mathscr{I}\colon\mathbb{H}\rightarrow\mathcal{H}$ exists such that $\mathscr{I}^{-1}Z\mathscr{I}=J_{0}$ then, as one may easily check, the set $\{\mathscr{I}e_{n}\}_{n\in\mathbb{N}}$ — where $\bm{e}\equiv\{e_{n}\}_{n\in\mathbb{N}}$ is the canonical basis in $\mathbb{H}$ — provides a $Z$-invariant orthonormal basis in $\mathcal{H}$. Conversely, assuming the existence of an orthonormal basis $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$ invariant under the action of $Z$, we have, for every $\chi$ in $\mathcal{H}$, that
$Z\chi=Z\Big(\sum_{n}\langle\psi_{n},\chi\rangle\psi_{n}\Big)=\sum_{n}\overline{\langle\psi_{n},\chi\rangle}Z\psi_{n}=\sum_{n}\overline{\langle\psi_{n},\chi\rangle}\psi_{n},$ (113)
and from this, it is clear that $\mathscr{I}\equiv W_{\Psi}^{-1}$ provides the isomorphism claimed in the proposition. Therefore, the proposition is proved if we can show that it is not always possible to construct a $Z$-invariant orthonormal basis in $\mathcal{H}$. To this end, let $\mathfrak{S}$ denote set of all the elements in $\mathcal{H}$ invariant under the action of $Z$, i.e.,
$\mathfrak{S}\coloneqq\{\eta\in\mathcal{H}\mid Z\eta=\eta\}.$ (114)
Note that $\mathfrak{S}$ is not empty, as it contains at least the null vector; moreover, although it is not linear on $\mathbb{Q}_{p,\mu}$, it is a $\mathbb{Q}_{p}$-linear space since, for any $\lambda_{1},\lambda_{2}\in\mathbb{Q}_{p}$ and $\eta_{1},\eta_{2}$ in $\mathfrak{S}$, we have that $\lambda_{1}\eta_{1}+\lambda_{2}\eta_{2}$ belongs to $\mathfrak{S}$ too. It is clear that $\mathfrak{S}$ is everywhere dense in $\mathcal{H}$: It suffices to observe that any $\chi$ in $\mathcal{H}$ can be expressed as
$\chi=\frac{1}{2}\big(\chi_{1}+\sqrt{\mu}\chi_{2}\big),$ (115)
where we set
$\chi_{1}\equiv Z\chi+\chi,\;\;\mbox{and}\;\;\chi_{2}=\frac{1}{\sqrt{\mu}}\big(Z\chi-\chi\big),$ (116)
with $Z\chi_{1}=\chi_{1}$ and $Z\chi_{2}=\chi_{2}$, i.e., $\chi_{1},\chi_{2}$ are $Z$-invariant. Thus, every $\chi$ in $\mathcal{H}$ can be expressed as a linear combination of elements in $\mathfrak{S}$. It then follows that a $Z$-invariant orthonormal basis in $\mathcal{H}$ exists if and only if we can construct an orthonormal basis in $\mathfrak{S}$. Let $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$ be an orthonormal basis in $\mathcal{H}$. It is clear that a $Z$-invariant dense subset $\Psi\equiv\{\psi_{n}\}_{n\in\mathbb{N}}$ in $\mathcal{H}$ — i.e., a dense subset in $\mathfrak{S}$ — can be constructed by setting
$\psi_{2n}\coloneqq Z\phi_{n}+\phi_{n},\quad\psi_{2n-1}\coloneqq\frac{1}{\sqrt{\mu}}\big(Z\phi_{n}-\phi_{n}\big).$ (117)
However, we know that this set is normal and IP-orthogonal, but cannot be made orthonormal without violating the $Z$-invariance condition (see Example 4.9). Hence, in general, an orthonormal basis in $\mathfrak{S}$ — i.e., a $Z$-invariant basis in $\mathcal{H}$ — cannot be constructed, which is what we want to prove. ∎
Remark 4.11.
A conjugation operator $J$, in a complex Hilbert space $\mathscr{K}$, is defined as an involutive anti-unitary operator. Then, it is not difficult to prove that there exists an orthonormal basis $\Psi\equiv\{\psi_{k}\}_{k\in\mathbb{N}}$ in $\mathscr{K}$ which is invariant under the action of $J$, i.e., such that $J\psi_{k}=\psi_{k}$, for all $k\in\mathbb{N}$ [28]. Indeed, let $\mathfrak{G}$ denote the set of all vectors $\chi$ in $\mathscr{K}$ invariant under the action of $J$, i.e., such that $J\chi=\chi$; this set is easily seen to be dense in $\mathscr{K}$. Since it is non-empty, it is separable and, thus, contains an everywhere dense subset. Applying to this set the so-called Gram–Schmidt orthonormalization process yields a complete orthonormal subset of $\mathfrak{G}$ and — since $\mathfrak{G}$ is dense — a $J$-invariant basis in $\mathscr{K}$. In the $p$-adic setting, what prevents us to repeat a similar general construction is the lack of a $p$-adic counterpart of the Gram–Schmidt process, which precludes the possibility to turn a normal set into an orthonormal basis in the invariant subspace — thus, an invariant orthonormal basis in the Hilbert space.
Remark 4.12.
Recall that, since $\mathbb{Q}_{p,\mu}$ is separable (and spherically complete [7]), every $p$-adic Hilbert space $\mathcal{H}$ is of countable-type — equivalently, separable. In fact, any subspace of $\mathcal{H}$ itself is of countable-type (cf. Corollary $2.3.14$ in [29]), and, thus, admits a $t$-orthogonal basis for some $t\in(0,1]$ (see Theorem $2.3.7$ in [29]). Owing on these observations, the conclusions of Proposition 4.10 are as expected. Specifically, if $Z$ is an involutive anti-unitary operator in $\mathcal{H}$, and $\mathfrak{S}$ a $Z$-invariant dense subspace, since $\mathfrak{S}$ is of countable-type, it always admits a $t$-orthogonal basis, for some $t\in(0,1]$. Then, since $\mathfrak{S}$ is dense in $\mathcal{H}$, this basis extends to a $Z$-invariant, $t$-orthogonal basis in the whole Hilbert space $\mathcal{H}$. Nevertheless, there is no guarantee that an orthonormal basis exists in $\mathfrak{S}$ — indeed, such a basis exists if and only if $\mathfrak{S}$ is an Hilbert subspace of $\mathcal{H}$ [23] — and, a fortiori, that a $Z$-invariant orthonormal basis can be constructed in the Hilbert space $\mathcal{H}$.
4.2 Construction of the isomorphism
Let $\Phi=\{\phi_{i}\}_{i\in\mathbb{N}}$ and $\Psi=\{\psi_{j}\}_{j\in\mathbb{N}}$ be two orthonormal bases in the $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$ respectively. Remember that $\{{{}^{ji}}E^{\Phi,\Psi}\}_{j,i\in\mathbb{N}}=\{\lvert\psi_{j}\rangle\langle\phi_{i}\rvert\}_{j,i\in\mathbb{N}}\equiv\{\langle\phi_{i},\cdot\rangle\psi_{j}\}_{j,i\in\mathbb{N}}$ is an orthonormal basis for $\mathcal{T}(\mathcal{H},\mathcal{K})$ (see Theorem 2.24). Therefore, any $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$ can be written, in Dirac notation, as
$T=\sum_{j,i\in\mathbb{N}}T_{ji}\lvert\psi_{j}\rangle\langle\phi_{i}\rvert,\quad(T_{ji})\in\mathsf{M}_{\infty}(\mathbb{Q}_{p,\mu}).$ (118)
For an operator $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$, we also have that
$\langle{{}^{ji}}E^{\Phi,\Psi},T\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\mathrm{tr}(|\psi_{j}\rangle\langle\phi_{i}|T)=T_{ji}.$ (119)
Proposition 4.13.
Let $J_{\Phi}$ be the conjugation operator associated with the orthonormal basis $\Phi\equiv\{\phi_{i}\}_{i\in\mathbb{N}}\subset\mathcal{H}$ (cf. Definition 4.2), and let $\Psi\equiv\{\psi_{j}\}_{j\in\mathbb{N}}$ be an orthonormal basis in $\mathcal{K}$. Let us define the map
$\mathcal{I}\colon\mathcal{T}(\mathcal{H},\mathcal{K})\ni T\mapsto\mathcal{I}(T)\coloneqq\sum_{i,j\in\mathbb{N}}\langle\psi_{j},T\phi_{i}\rangle(\phi_{i}\otimes\psi_{j})\in\mathcal{H}\otimes\mathcal{K}.$ (120)
Then, we have that
$\mathcal{I}(\lvert w\rangle\langle v\rvert)=(J_{\Phi}v)\otimes w$ (121)
holds for every vector $v=\sum_{i\in\mathbb{N}}\langle\phi_{i},v\rangle\phi_{i}$ in $\mathcal{H}$ and $w=\sum_{j\in\mathbb{N}}\langle\psi_{j},v\rangle\psi_{j}$ in $\mathcal{K}$, such that $\lvert w\rangle\langle v\rvert\in\mathcal{T}(\mathcal{H},\mathcal{K})$.
Proof.
The following chain of equations holds
[…]
$=\sum_{i,j\in\mathbb{N}}\langle v,\phi_{i}\rangle\langle\psi_{j},w\rangle(\phi_{i}\otimes\psi_{j})$
$=\bigg(\sum_{i\in\mathbb{N}}\langle v,\phi_{i}\rangle\phi_{i}\bigg)\otimes\bigg(\sum_{j\in\mathbb{N}}\langle\psi_{j},w\rangle\psi_{j}\bigg)=(J_{\Phi}v)\otimes w.$ (122)
Next result provides the desired isomorphism between the tensor product of $p$-adic Hilbert spaces $\mathcal{H}\otimes\mathcal{K}$, and the Hilbert space of $p$-adic Hilbert-Schmidt operators $\mathcal{T}(\mathcal{H},\mathcal{K})$.
Theorem 4.14.
The application $\mathcal{I}$ defined in Proposition 4.13 is an inner-product-preserving linear isomorphism between the $p$-adic Hilbert spaces $\mathcal{H}\otimes\mathcal{K}$ and $\mathcal{T}(\mathcal{H},\mathcal{K})$. Moreover, the map
$\mathcal{I^{*}}\colon\mathcal{H}\otimes\mathcal{K}\ni u\mapsto\mathcal{I^{*}}(u)\coloneqq\sum_{i,j\in\mathbb{N}}\langle\phi_{i}\otimes\psi_{j},u\rangle\lvert\psi_{j}\rangle\langle\phi_{i}\rvert\in\mathcal{T}(\mathcal{H},\mathcal{K})$ (123)
is such that $\mathcal{I}^{*}=\mathcal{I}^{-1}.$
Proof.
Since $\mathcal{I}$ is linear and maps the norm-orthogonal basis $\{\lvert\psi_{j}\rangle\langle\phi_{i}\rvert\}_{j,i\in\mathbb{N}}$ to $\{(J_{\Phi}\phi_{i})\otimes\psi_{j}\}_{i,j\in\mathbb{N}},$ then it is inner-product preserving and injective. To prove it is also surjective, let us take $u\in\mathcal{H}\otimes\mathcal{K}$, and let $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$ be written as
$T=\sum_{i,j\in\mathbb{N}}\langle\phi_{j}\otimes\psi_{i},u\rangle\lvert\psi_{i}\rangle\langle\phi_{j}\rvert$. (124)
Then, we have that
$\begin{split}\mathcal{I}(T)&\coloneqq\sum_{i,j\in\mathbb{N}}\langle\psi_{j},T\phi_{i}\rangle(\phi_{i}\otimes\psi_{j})=\sum_{i,j\in\mathbb{N}}\langle\psi_{j},\sum_{s,t\in\mathbb{N}}\langle\phi_{s}\otimes\psi_{t},u\rangle\psi_{t}\rangle\langle\phi_{s},\phi_{i}\rangle(\phi_{i}\otimes\psi_{j})\\ &\>=\sum_{s,t\in\mathbb{N}}\langle\phi_{s}\otimes\psi_{t},u\rangle\sum_{i,j\in\mathbb{N}}\delta_{j,t}\delta_{s,i}(\phi_{i}\otimes\psi_{j})=\sum_{i,j\in\mathbb{N}}\langle\phi_{i}\otimes\psi_{j},u\rangle(\phi_{i}\otimes\psi_{j})=u.\end{split}$ (125)
By Corollary $6.13$ in [22], for an operator $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$ we have $\lim_{i+j}\langle\psi_{j},T\phi_{i}\rangle=0.$ This, together with Remark $6.5$ in [22], entails that the double series
$\sum_{i\in\mathbb{N}}\sum_{j\in\mathbb{N}}\langle\psi_{j},T\phi_{i}\rangle=\sum_{j\in\mathbb{N}}\sum_{i\in\mathbb{N}}\langle\psi_{j},T\phi_{i}\rangle=\sum_{i,j\in\mathbb{N}}\langle\psi_{j},T\phi_{i}\rangle$ (126)
converges, and we can exchange the order of the sums in the second equality of (125). Since (125) holds for all $u\in\mathcal{H}\otimes\mathcal{K}$, the application $\mathcal{I}$ is surjective. Moreover, by observing that
$\langle\mathcal{I}^{*}\mathcal{I}(T),S\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\langle\mathcal{I}(T),\mathcal{I}(S)\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\langle T,S\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}},$ (127)
we can argue that $\mathcal{I}^{\ast}\mathcal{I}=\mathrm{Id}$ and $\mathcal{I}\mathcal{I}^{\ast}=\mathrm{Id}$, i.e., $\mathcal{I}^{\ast}=\mathcal{I}^{-1}$.
To prove (123), it suffices to note that a straightforward computation yields
$\bigg\langle T,\sum_{i,j\in\mathbb{N}}\langle\phi_{i}\otimes\psi_{j},u\rangle\lvert\psi_{j}\rangle\langle\phi_{i}\rvert\bigg\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\langle\mathcal{I}(T),u\rangle=\langle T,\mathcal{I}^{*}(u)\rangle.$ (128)
for every $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$. ∎
Corollary 4.15.
For the operator $\lvert w\rangle\langle v\rvert\in\mathcal{T}(\mathcal{H},\mathcal{K})$ — where $v\in\mathcal{H}$ and $w\in\mathcal{K}$ are any two vectors — the following equalities hold
  1. (i)
    $J_{\Phi}v=v$;
  2. (ii)
    $\lvert w\rangle\langle v\rvert=\mathcal{I}^{*}((J_{\Phi}v)\otimes w)$;
  3. (iii)
    $\mathcal{I}^{*}(v\otimes w)=|w\rangle\langle J_{\Phi}v|$.
Remark 4.16.
We observe that the relation between the $p$-adic Hilbert-Schmidt-class operators $\mathcal{T}(\mathcal{H},\mathcal{K})$ and the tensor product of $p$-adic Hilbert spaces $\mathcal{H}\otimes\mathcal{K}$ is essentially related to the correspondence $\lvert\psi\rangle\langle\phi\rvert\leftrightarrow\phi\otimes\psi,$ with necessary conjugation operators, as it happens in the complex setting.
5 Tensor product of subspaces
In this section, we address the tensor product structure of subspaces in a (tensor product) $p$-adic Hilbert space. Once again, the well known results from the complex framework [26] will provide us a useful road map, but, when switching to the $p$-adic setting, the emergence of non-trivial and distinctive peculiarities are encountered.
In what follows, for the sake of notational convenience, we set $c_{0}(I,\mathbb{Q}_{p,\mu})\equiv c_{0}(I)$ — where $I$ is a finite set, or $I=\mathbb{N}$. We start by proving
Proposition 5.1.
Let $(X,\|\cdot\|)$ be a $p$-adic Banach space. Then, there exists a canonical $p$-adic Banach space isomorphism between $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ and $c_{0}(I,X)$.
Proof.
To construct the isomorphism, we proceed in two steps: First, we prove that there exists a natural isomorphism between $c_{0}(I)\otimes_{\pi}X$ and $c_{0}(I,X)$; then, we show that this map admits a unique extension to the complete tensor product space $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$. Specifically, let us consider the map $\mathcal{L}_{X}\colon c_{0}(I)\otimes_{\pi}X\rightarrow c_{0}(I,X)$ defined by
$c_{0}(I)\otimes_{\pi}X\ni\lambda\otimes x\mapsto\mathcal{L}_{X}(\lambda\otimes x)\coloneqq\{\lambda_{i}x\}_{i\in I}\in c_{0}(I,X),$ (129)
where $\lambda=\{\lambda_{i}\}_{i\in I}\in c_{0}(I)$, and $x\in X$. Since $\lambda\in c_{0}(I)$, then, for every $x$ in $X$, it is clear that $\{\lambda_{i}x\}_{i\in I}$ is an element of $c_{0}(I,X)$, i.e.,
$\lim_{i}\lambda_{i}x=0,\quad\|\{\lambda_{i}x\}_{i\in I}\|_{\infty}=\sup_{i}\left|\lambda_{i}\right|\,\|x\|=\|\lambda\|_{\infty}\,\|x\|.$ (130)
We now aim to show that $\mathcal{L}_{X}$ defines a norm-preserving map between $c_{0}(I)\otimes_{\pi}X$ and $c_{0}(I,X)$. Indeed, let $u=\sum_{j=1}^{m}\lambda_{j}\otimes x_{j}$ be a possible representation of $u\in c_{0}(I)\otimes_{\pi}X$ where, without loss of generality, we can assume $\{x_{j}\}_{j=1}^{m}$ to be a norm-orthogonal system in $X$. Taking into account that, for $j=1,\ldots,m$, $\lambda_{j}=\{\lambda_{i,j}\}_{i\in I}$, we have:
$\mathcal{L}_{X}(u)\equiv\{u_{i}\}_{i\in I}=\bigg\{\sum_{j=1}^{m}\lambda_{i,j}x_{j}\bigg\}_{i\in I}.$ (131)
Therefore, it follows that
$\|\mathcal{L}_{X}(u)\|_{\infty}=\bigg\|{\bigg\{\sum_{j=1}^{m}\lambda_{i,j}x_{j}\bigg\}}_{i\in I}\bigg\|_{\infty}=\sup_{i\in I}\bigg\|\sum_{j=1}^{m}\lambda_{i,j}x_{j}\bigg\|$ $\leq\sup_{i\in I}\max_{1\leq j\leq m}\left|\lambda_{i,j}\right|\,\|x_{j}\|$
$=\max_{1\leq j\leq m}\|\lambda_{j}\|_{\infty}\,\|x_{j}\|=\|u\|_{\pi}.$ (132)
To prove the reverse inequality, let us fix a representation $\sum_{j=1}^{m}\lambda_{j}\otimes x_{j}$ of $u$ in $c_{0}(I)\otimes_{\pi}X$ — where, again, we assume $\{x_{j}\}_{j=1}^{m}$ to be a norm-orthogonal system in $X$. Let us consider the series $\sum_{n\in I}e_{n}\otimes u_{n}$, where $\{u_{n}\}_{n\in I}=\{\sum_{j=1}^{m}\lambda_{n,j}x_{j}\}_{n\in I}=\mathcal{L}(u)$, and where $\bm{e}=\{e_{n}\}_{n\in I}$ is the so-called canonical basis in $c_{0}(I)$ (cf. Example 2.11). Let $\Pi_{k}\colon c_{0}(I)\rightarrow c_{0}(K)$ — where $K=\{1,\ldots,k\}$, $\mathrm{card}(K)\leq\mathrm{card}(I)$ — denote the projection of the elements in $c_{0}(I)$ onto the first $k$ coordinates. We then have that
$\Pi_{k}(\lambda)=\sum_{n=1}^{k}\lambda_{n}e_{n},\quad\forall\lambda=\{\lambda_{n}\}_{n\in I}\in c_{0}(I).$ (133)
It is clear that $\Pi_{k}(\lambda)\rightarrow\lambda$ for $k\rightarrow\infty$ (if $I$ is a finite set, for $k$ equal to $\text{card}(I)$), and, therefore,
$\bigg\|u-\sum_{n=1}^{k}e_{n}\otimes u_{n}\bigg\|_{\pi}$ $=\bigg\|\sum_{j=1}^{m}\lambda_{j}\otimes x_{j}-\sum_{n=1}^{k}\sum_{j=1}^{m}e_{n}\otimes\lambda_{n,j}x_{j}\bigg\|_{\pi}$
$=\bigg\|\sum_{j=1}^{m}\bigg(\lambda_{j}\otimes x_{j}-\sum_{n=1}^{k}\lambda_{n,j}e_{n}\otimes x_{j}\bigg)\bigg\|_{\pi}$
$=\bigg\|\sum_{j=1}^{m}\big(\lambda_{j}-\Pi_{k}(\lambda_{j})\big)\otimes x_{j}\bigg\|_{\pi}$
$\leq\max_{1\leq j\leq m}\|\lambda_{j}-\Pi_{k}(\lambda_{j})\|_{\infty}\,\|x_{j}\|.$ (134)
Thus, for $k\rightarrow\infty$ (or, if $I$ is a finite set, for $k$ equal to $\text{card}(I)$), $\|u-\sum_{n=1}^{k}e_{n}\otimes u_{n}\|_{\pi}=0$, and, hence, we proved that $\sum_{n=1}^{k}e_{n}\otimes u_{n}$ converges to $u$ in the projective norm $\|\cdot\|_{\pi}$. Now, we have:
$\|u\|_{\pi}=\bigg\|\sum_{i\in I}e_{i}\otimes u_{i}\bigg\|_{\pi}\leq\sup_{i}\|u_{i}\|\,\|e_{i}\|_{\infty}=\sup_{i}\|u_{i}\|=\|\mathcal{L}_{X}(u)\|_{\infty}.$ (135)
Therefore, taking into account (5) and (135), we conclude that $\|\mathcal{L}_{X}(u)\|_{\infty}=\|u\|_{\pi}$, and, hence, that $\mathcal{L}_{X}$ is a norm-preserving map (an isometry) between the $p$-adic normed spaces $c_{0}(I)\otimes_{\pi}X$ and $c_{0}(I,X)$. On the other hand, since $c_{0}(I,X)$ is complete, $\mathcal{L}_{X}$ admits a unique bounded (Hahn-Banach) extension to an isometric operator — which we still denote with the same symbol — from the completed projective tensor product $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ into $c_{0}(I,X)$.
To prove that $\mathcal{L}_{X}$ is an isomorphism of $p$-adic Banach spaces, we are left to show that it is a surjective map. To this end, let $x=\{x_{i}\}_{i\in I}$ be a sequence in $c_{0}(I,X)$. If we consider the series $\sum_{i\in I}e_{i}\otimes x_{i}$, it is clear that its image under $\mathcal{L}_{X}$ coincides with $x=\{x_{i}\}_{i\in I}$. Hence, we have only to show that $\sum_{i\in I}e_{i}\otimes x_{i}$ converges in $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$. If $\mathrm{card}(I)<\infty$, this is certainly true. On the other hand, if $I=\mathbb{N}$, since $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ is complete, it suffices to show that $\big\{\sum_{i=1}^{n}e_{i}\otimes x_{i}\big\}_{n\in\mathbb{N}}$ is a $p$-adic Cauchy sequence. This simply follows by observing that [^1] We recall that, if $(X,\|\cdot\|)$ is a $p$-adic normed space, a sequence $\{x_{n}\}_{n\in\mathbb{N}}$ is a Cauchy sequence in $X$ iff $\lim_{n}\|x_{n+1}-x_{n}\|=0$. See Chapter $1$ in [30].
$\bigg\|\sum_{i=1}^{n+1}e_{i}\otimes x_{i}-\sum_{i=1}^{n}e_{i}\otimes x_{i}\bigg\|_{\pi}=\|e_{n+1}\otimes x_{n+1}\|_{\pi}=\|e_{n+1}\|_{\infty}\,\|x_{n+1}\|,$ (136)
and recalling that, since $\{x_{n}\}_{n\in I}\in c_{0}(I,X)$, then $\lim_{n}x_{n}=0$. Hence, $\mathcal{L}_{X}$ is a surjective isometry — i.e., an isometric isomorphism — between the $p$-adic Banach spaces $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ and $c_{0}(I,X)$. ∎
Proposition 5.1 allows us to identify the tensor product $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ with the $p$-adic Banach space $c_{0}(I,X)$ of zero-convergent sequences in $X$. As a primary implication of this result, we now prove the following
Corollary 5.2.
Let $(Y,\|\cdot\|)$ be a $p$-adic Banach space, and let $X\subset Y$ be a closed subspace of $Y$. Then, $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ is a subspace of $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}Y$.
Proof.
From Proposition 5.1, we know that the map $\mathcal{L}_{Y}\colon c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}Y\rightarrow c_{0}(I,Y)$ (respectively $\mathcal{L}_{X}\colon c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X\rightarrow c_{0}(I,X)$) provides an isometric isomorphism of $p$-adic Banach spaces; i.e.,
$c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}Y\cong c_{0}(I,Y),\quad c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X\cong c_{0}(I,X).$ (137)
On the other hand, $c_{0}(I,X)$ is a subspace of $c_{0}(I,Y)$ and, hence, $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}X$ is a subspace of $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}Y$, that is what we wanted to prove. ∎
Remark 5.3.
Proposition 5.2 is a $p$-adic counterpart of a well known result for the projective tensor product of complex Banach spaces. Specifically, let $\mathscr{X}$ and $\mathscr{Y}$ be two (complex) Banach spaces, and let $\mathscr{W}$ be a closed subspace of $\mathscr{Y}$. Denoting by $\mathscr{X}\operatorname{\widehat{\otimes}_{\pi}}\mathscr{W}$ the completion of the projective tensor product of $\mathscr{X}$ and $\mathscr{W}$, it is readily seen that $\mathscr{X}\operatorname{\widehat{\otimes}_{\pi}}\mathscr{W}$ will not be, in general, a subspace of $\mathscr{X}\operatorname{\widehat{\otimes}_{\pi}}\mathscr{Y}$, unless $\mathscr{X}=\ell_{1}(I)$ is the Banach space of absolutely convergent complex sequences indexed by $I$ — in this case, one can indeed show that $\ell_{1}(I)\operatorname{\widehat{\otimes}_{\pi}}\mathscr{W}$ is a subspace of $\ell_{1}(I)\operatorname{\widehat{\otimes}_{\pi}}\mathscr{Y}$; see Proposition $2.7$ in [26]. Therefore, the result of Proposition 5.2 is as expected once observed that, in the $p$-adic setting, $\ell_{1}(I)\equiv c_{0}(I)$ [39].
Our aim is now to investigate how the tensor product operation behaves w.r.t. the $p$-adic Hilbert subspace structure. We start by recalling that, in the $p$-adic setting, there are two natural notions of subspaces that can be defined for a $p$-adic Hilbert space (see [23] for a detailed discussion); specifically, we have:
Definition 5.4.
Let $\mathcal{H}$ be a $p$-adic Hilbert space, and let $\mathcal{W}\subset\mathcal{H}$ be a norm-closed subspace of $\mathcal{H}$. We say that $\mathcal{W}$ is a Hilbert subspace of $\mathcal{H}$ if it is a $p$-adic Hilbert space itself — equivalently, if $\mathcal{W}$ admits an orthonormal basis $\Psi\equiv\{\psi_{i}\}_{i\in I}\subset\mathcal{W}$, $\mathrm{card}(I)=\dim\mathcal{W}$. We say that $\mathcal{W}$ is a regular subspace of $\mathcal{H}$, if it is a Hilbert subspace, and if some (hence, any) orthonormal basis of $\mathcal{W}$ can be extended to an orthonormal basis of $\mathcal{H}$.
Clearly, a regular subspace is a Hilbert subspace, while the contrary is not generally true.
Remark 5.5.
Let $\mathcal{W}$ be a norm-closed subspace of a $p$-adic Hilbert space $\mathcal{H}$. Denoting by $\mathcal{W}^{\perp}$ the (closed) subspace defined as
$\mathcal{W}^{\perp}\coloneqq\{x\in\mathcal{H}\mid\langle x,w\rangle=0,\;\;\forall w\in\mathcal{W}\},$ (138)
it is possible to prove that if $\mathcal{W}$ is a regular subspace, then $\mathcal{W}^{\perp}$ is a regular subspace too. If $\mathcal{W}^{\perp}$ is a Hilbert subspace of $\mathcal{H}$, and if the relation
$\|w+v\|=\max\{\|w\|,\|v\|\}$ (139)
holds for every $v\in\mathcal{W}$ and $w\in\mathcal{W}^{\perp}$, then $\mathcal{W}$ and $\mathcal{W}^{\perp}$ are regular subspaces of $\mathcal{H}$ [23].
Proposition 5.6.
Let $\mathcal{H}$ and $\mathcal{K}$ be two $p$-adic Hilbert spaces, and let $\Psi\equiv\{\psi_{i}\}_{i\in I}$ be an orthonormal basis in $\mathcal{H}$. Then, the map defined as
$\mathscr{S}_{\Psi}\colon\mathcal{H}\otimes\mathcal{K}\ni u\mapsto\mathscr{S}_{\Psi}(u)\coloneqq\sum_{j=1}^{m}W_{\Psi}(x_{j})\otimes y_{j}\in\mathbb{H}(I)\otimes\mathcal{K}$ (140)
— where $W_{\Psi}\colon\mathcal{H}\rightarrow\mathbb{H}(I)$ is the canonical isomorphism of $\mathcal{H}$ into the coordinate Hilbert space $\mathbb{H}(I)$, and $u=\sum_{j=1}^{m}x_{j}\otimes y_{j}$, $x_{j}\in\mathcal{H}$, $y_{j}\in\mathcal{K}$, any representation of $u$ — is an isomorphism between the $p$-adic Hilbert spaces $\mathcal{H}\otimes\mathcal{K}$ and $\mathbb{H}(I)\otimes\mathcal{K}$.
Proof.
It is clear that $\mathscr{S}_{\Psi}$ is a surjective map. Moreover, by observing that
$\|\mathscr{S}_{\Psi}(u)\|_{\pi}=\bigg\|\sum_{j=1}^{m}W_{\Psi}(x_{j})\otimes y_{j}\bigg\|$ $=\inf\bigg\{\max_{1\leq j\leq m}\|W_{\Psi}(x_{j})\|\,\|y_{j}\|\,\,\big\lvert\,\,u=\sum_{j=1}^{m}x_{j}\otimes y_{j},\,\,x_{j}\in\mathcal{H},y_{j}\in\mathcal{K}\bigg\}$
$=\inf\bigg\{\max_{1\leq j\leq m}\|x_{j}\|\,\|y_{j}\|\,\,\big\lvert\,\,u=\sum_{j=1}^{m}x_{j}\otimes y_{j}\,\,x_{j}\in\mathcal{H},y_{j}\in\mathcal{K}\bigg\}$
$=\|u\|_{\pi},$ (141)
we can argue that $\mathscr{S}_{\Psi}$ is norm-preserving. Next, let $u=\sum_{j=1}^{m}x_{j}\otimes y_{j}$ and $w=\sum_{k=1}^{n}f_{k}\otimes g_{k}$, where $x_{j},f_{k}\in\mathcal{H}$, $y_{j},g_{k}\in\mathcal{K}$, be two elements of $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$. Then, we have:
$\langle\mathscr{S}_{\Psi}(u),\mathscr{S}_{\Psi}(w)\rangle_{\pi}$ $=\Big\langle\sum_{j=1}^{m}W_{\Psi}(x_{j})\otimes y_{j},\sum_{k=1}^{n}W_{\Psi}(f_{k})\otimes g_{k}\Big\rangle_{\pi}$
$=\sum_{j=1}^{m}\sum_{k=1}^{n}\big\langle W_{\Psi}(x_{j}),W_{\Psi}(f_{k})\big\rangle_{\mathcal{H}}\langle y_{j},g_{k}\rangle_{\mathcal{K}}$
$=\sum_{j=1}^{m}\sum_{k=1}^{n}\langle\breve{x}_{j},\breve{f}_{k}\rangle_{\mathbb{H}(I)}\langle y_{j},g_{k}\rangle_{\mathcal{K}}.$ (142)
[…]
We now come to the main result of this section
Proposition 5.7.
Let $\mathcal{H}$, $\mathcal{K}$ be two $p$-adic Hilbert spaces, and let $\mathcal{W}$ be a Hilbert subspace of $\mathcal{K}$. Then $\mathcal{H}\otimes\mathcal{W}$ is a Hilbert subspace of $\mathcal{H}\otimes\mathcal{K}$. Moreover, if $\mathcal{W}$ is regular, then $\mathcal{H}\otimes\mathcal{W}$ is a regular subspace of $\mathcal{H}\otimes\mathcal{K}$.
Proof.
Let $\Psi\equiv\{\psi_{i}\}_{i\in I}$ be an orthonormal basis in $\mathcal{H}$ ($\dim\mathcal{H}=\mathrm{card}(I)$). The map $\mathscr{S}_{\Psi}$ in Proposition 5.6 provides an isomorphism between the $p$-adic Hilbert spaces $\mathcal{H}\otimes\mathcal{K}$ and $\mathbb{H}(I)\otimes\mathcal{K}$. In particular, this also entails that $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}\cong c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$ are isomorphic $p$-adic Banach spaces (recall that, we denoted with $c_{0}(I)$ the carrier Banach space of $\mathbb{H}(I)$, see Example 2.11). In a similar way, if $\mathcal{W}$ is a Hilbert subspace of $\mathcal{K}$, we have the ($p$-adic Banach space isomorphism) $\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{W}\cong c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}\mathcal{W}$. By Proposition 5.1, we know that $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}\mathcal{W}$ is a subspace of $c_{0}(I)\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K}$ — and, therefore, that $\mathbb{H}(I)\otimes\mathcal{W}$ is a subspace of $\mathbb{H}(I)\otimes\mathcal{K}$, as well. Then, we have the following inclusion
$\mathcal{H}\otimes\mathcal{W}\cong\mathbb{H}(I)\otimes\mathcal{W}\subset\mathbb{H}(I)\otimes\mathcal{K}\cong\mathcal{H}\otimes\mathcal{K},$ (143)
and since $\mathcal{H}\otimes\mathcal{W}$ is a Hilbert space itself — and norm-closed in $\mathcal{H}\otimes\mathcal{K}$ — it is a Hilbert subspace of $\mathcal{H}\otimes\mathcal{K}$. Assume now $\mathcal{W}\subset\mathcal{K}$ be a regular subspace. This entails that there exists an orthonormal basis $\Phi_{K}\equiv\{\phi_{k}\}_{k\in K}$ of $\mathcal{K}$ such that, for some subset $J$ of $K$ — putting $\Phi_{J}\equiv\{\phi_{j}\}_{j\in J}$$\Phi_{J}$ is an orthonormal basis of $\mathcal{W}$. Clearly, $\mathcal{W}$ is a Hilbert subspace of $\mathcal{K}$; thus, $\mathcal{H}\otimes\mathcal{W}$ is a Hilbert subspace of $\mathcal{H}\otimes\mathcal{K}$. Moreover, from Theorem 3.10, we know that the set $\{\psi_{i}\otimes\phi_{j}\}_{i\in I,j\in J}$ — where $\Psi_{I}\equiv\{\psi_{i}\}_{i\in I}$ is an orthonormal basis in $\mathcal{H}$ — provides an orthonormal basis in $\mathcal{H}\otimes\mathcal{W}$. Then, the set $\{\psi_{i}\otimes\phi_{k}\}_{i\in I,k\in K}$ — obtained by extending the basis $\Phi_{J}\subset\mathcal{W}$ to the basis $\Phi_{K}$ of $\mathcal{K}$ — is, in a natural way, an orthonormal basis of $\mathcal{H}\otimes\mathcal{K}$. Concluding, we have constructed an orthonormal basis in $\mathcal{H}\otimes\mathcal{W}$ which can be extended to an orthonormal basis of $\mathcal{H}\otimes\mathcal{K}$, i.e., $\mathcal{H}\otimes\mathcal{W}$ is a regular Hilbert subspace of $\mathcal{H}\otimes\mathcal{K}$. ∎
6 Conclusions
In this work we have introduced the tensor product of $p$-adic Hilbert spaces as a fundamental step in advancing the abstract approach theory of quantum mechanics on a quadratic extension $\mathbb{Q}_{p,\mu}$ of $p$-adic numbers [24, 22, 23]. Let us briefly summarize our main achievements:
  • We defined the algebraic tensor product $\mathcal{H}\otimes_{\alpha}\mathcal{K}$ between $p$-adic Hilbert spaces considering them as simple vector spaces, i.e. without additional structure. Then, the key point was the introduction of the norm $\|\cdot\|_{\pi},$ which is the most natural adaptation, to our case of study, of the projective norm used in tensor product of complex Banach spaces. Metrically completing the $p$-adic normed space obtained (after providing a suitable inner product $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi}$ together with an orthonormal basis $\bm{e}$), we finally obtained the tensor product of $p$-adic Hilbert spaces $\Big(\mathcal{H}\operatorname{\widehat{\otimes}_{\pi}}\mathcal{K},\|\cdot\|_{\pi},\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\pi},\bm{e}\equiv\{e_{k}\}_{k\in K}\equiv\{\phi_{i}\otimes\psi_{j}\}_{i\in I,\,j\in J}\Big)$.
  • We characterized the space of Hilbert-Schmidt-class operators $\mathcal{T}(\mathcal{H},\mathcal{K})$ between any two $p$-adic Hilbert spaces which we know is equivalent to the space of $p$-adic trace class operators between $\mathcal{H}$ and $\mathcal{K}$.
  • We defined the conjugation operator $J_{\Phi}$ and then the set of anti-unitary operators $\overline{\mathcal{U}}(\mathcal{H})$ in a $p$-adic Hilbert space $\mathcal{H}$ as conjugations of unitary operators $\mathcal{U}(\mathcal{H})$. We used them to construct the isomorphism, similar to the one in the complex case, between the tensor product $\mathcal{H}\otimes\mathcal{K}$ and the space of trace class operators $\mathcal{T}(\mathcal{H},\mathcal{K}).$
  • We studied the tensor product of subspaces of a $p$-adic Hilbert space. We adapted some known results in the case of complex Banach spaces to our non-Archimedean case and, at the same time, we also analyzed the substantial differences. In particular, we focused on the tensor product of the $p$-adic Banach space of zero-converging sequences $c_{0}(I)$.
We left for future work the construction of the dual $(\mathcal{H}\otimes\mathcal{K})^{\prime}$ of $\mathcal{H}\otimes\mathcal{K}$, and the study of its properties, like whether or not there exists an isomorphism between $(\mathcal{H}\otimes\mathcal{K})^{\prime}$ and $\mathcal{H}\otimes\mathcal{K}$. Similarly, we did not address the construction of the bidual of $\mathcal{H}\otimes\mathcal{K}$ and, hence, the study of the reflexivity property of $\mathcal{H}\otimes\mathcal{K}$. Nevertheless, the present work paves the way for exploring the notion of entanglement in the $p$-adic setting. This can be pursued by a thorough study of the trace-class operators in $\mathcal{H}\otimes\mathcal{K}$.
Appendix A Proofs of Theorems 2.162.18 and 2.24
Here we provide detailed proofs of the results stated in Subsection 2.2. Since these theorems are proved likewise the analogous results for the single Hilbert space setting (see, in particular, Theorems 4.24.24.44.4 and $6.38$ in [22]), and only require minor modifications, we opted to place them in this appendix, so to leave the main text to the more original material. We start from
Proof of Theorem 2.16.
Let $\mathcal{H}$ and $\mathcal{K}$ be two $p$-adic Hilbert spaces with orthonormal bases $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$, and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ respectively. From relations (22), (23), it is clear that $\mathcal{B}(\mathcal{H},\mathcal{K})\subset(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$. To prove the reverse inclusion, and, thus, the first equality in (25), we have only to observe that if $B\equiv\mathrm{op}_{\Phi,\Psi}(B_{mn})\in(\mathcal{H},\mathcal{K})_{\Phi,\Psi}$ is an all-over matrix operator then, for any $\zeta$ in $\mathcal{H}$, we have:
$\|B\zeta\|=\sup_{m}\bigg|\sum_{n}B_{mn}\langle\phi_{n},\zeta\rangle\bigg|\leq\sup_{m,n}\left|B_{mn}\right|\,\left|\langle\phi_{n},\zeta\rangle\right|\leq\|\zeta\|\sup_{m,n}\left|B_{mn}\right|,$ (144)
i.e., $B\equiv\mathrm{op}_{\Phi,\Psi}(B_{mn})\in\mathcal{B}(\mathcal{H},\mathcal{K})$. The second equality in (25) simply follows by relations (22). To prove (26), one can notice that if $B$ is a bounded operator, then
$\|B\phi_{n}\|=\bigg\|\bigg(\sum_{m}\sum_{s}\langle\psi_{m},B\phi_{s}\rangle\langle\phi_{s},\cdot\rangle\psi_{m}\bigg)(\phi_{n})\bigg\|=\bigg\|\sum_{m}\sum_{s}\langle\psi_{m},B\phi_{s}\rangle\delta_{sn}\psi_{m}\bigg\|=\sup_{m}\left|B_{mn}\right|,$ (145)
where, in the first equality, we used the series expansion in (24) of the bounded operator $B$ — converging in the strong operator topology — while the last quality follows from the Parseval identity (cf. (11)). Hence, from (144), and taking into account relation (145), we can conclude that $\|B\|=\sup_{m,n}\left|B_{mn}\right|=\sup_{n}\|B\phi_{n}\|$. ∎
We next prove the main result providing the characterization of adjointable operators acting between two $p$-adic Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$.
Proof of Theorem 2.18.
Let $B$ be a bounded operator. Owing to the second series expansion in (24) we have, for all $\zeta\in\mathcal{K}$ and $\chi\in\mathcal{H}$, that
$\langle\zeta,B\chi\rangle=\sum_{n}\sum_{m}B_{mn}\langle\zeta,\psi_{m}\rangle\langle\phi_{n},\chi\rangle=\sum_{n}\bigg(\overline{\sum_{m}\overline{B_{mn}}\langle\psi_{m},\zeta\rangle}\bigg)\langle\phi_{n},\chi\rangle.$ (146)
Since $\chi$ is an arbitrary vector in $\mathcal{H}$, we can argue that
$\{\xi^{\zeta}_{n}\equiv\sum_{m}\overline{B_{mn}}\langle\psi_{m},\zeta\rangle\}_{n\in\mathbb{N}}\in\ell^{\infty},$ (147)
i.e., $\{\xi_{n}^{\zeta}\}_{n\in\mathbb{N}}$ is a bounded sequence. Using the isomorphism $\mathcal{L}_{\Phi}\colon\ell^{\infty}\rightarrow\mathcal{H}^{\prime}$ (see Remark 2.14), we have that
$\langle\zeta,B\chi\rangle=\sum_{n}\sum_{m}\overline{\xi_{n}^{\zeta}}\langle\phi_{n},\chi\rangle=\big(\mathcal{L}_{\Phi}(\overline{\xi^{\zeta}})\big)(\chi).$ (148)
Assume now that $B\in\mathcal{B}_{\mathrm{ad}}(\mathcal{H},\mathcal{K})$. Then, for some $\eta=B^{\ast}\zeta\in\mathcal{H}$, the following relation holds
$\big(\mathcal{L}_{\Phi}(\overline{\xi^{\zeta}})\big)(\chi)=\langle\zeta,B\chi\rangle=\langle B^{\ast}\zeta,\chi\rangle,\quad\forall\chi\in\mathcal{H};$ (149)
from this, and exploiting Proposition $3.36$ in [22], we can argue that $\{\xi_{n}^{\zeta}\}_{n\in\mathbb{N}}\in c_{0}$ and, moreover, that
$\sum_{n}\xi_{n}^{\zeta}\phi_{n}=B^{\ast}\zeta\in\mathcal{H}.$ (150)
Using the explicit form of $\xi_{n}^{\zeta}$ in (147), we can retrieve the matrix representation of the operator $B^{\ast}$, i.e.,
$B^{\ast}=\mathrm{op}_{\Psi,\Phi}(B^{\ast}_{mn})=\mathrm{op}_{\Psi,\Phi}(\overline{B_{nm}}).$ (151)
Concluding, the adjoint $B^{\ast}$ of $B$ is a bounded (see (31)), all-over matrix operator and, in the light of Theorem 2.16, its matrix coefficients $B^{\ast}_{mn}=\overline{B_{nm}}$ have to satisfy conditions (1)(3). By further taking into account that $\sup_{m,n}\left|\overline{B_{nm}}\right|=\sup_{m,n}\left|B_{mn}\right|$, relation (33) is also proved.
Conversely, let us suppose that $B=\mathrm{op}_{\Phi,\Psi}(B_{mn})$ is such that its matrix elements $B_{mn}$ satisfy conditions (1)-(3). Then, from (1) and (2), it follows that $B$ is an all-over bounded matrix operator, and expansion (146) holds for every $\zeta$ in $\mathcal{K}$ and $\chi$ in $\mathcal{H}$. Moreover if, for every $m\in\mathbb{N}$, $\lim_{n}B_{mn}=0$ then, setting $B^{\ast}=\mathrm{op}_{\Psi,\Phi}(B^{\ast}_{mn})=\mathrm{op}_{\Psi,\Phi}(\overline{B_{nm}})$, we have that the series $\sum_{m}\sum_{n}B^{\ast}_{mn}\langle\psi_{n},\zeta\rangle\phi_{m}$ must converge to some vector $\eta=\mathrm{op}_{\Psi,\Phi}(B_{mn}^{\ast})\zeta\in\mathcal{H}$. Therefore, for every $\zeta$ in $\mathcal{K}$, there is some $\eta=\mathrm{op}_{\Psi,\Phi}(B_{mn}^{\ast})\zeta$ in $\mathcal{H}$, such that
$\langle\zeta,B\chi\rangle=\langle B^{\ast}\zeta,\chi\rangle=\langle\eta,\chi\rangle.$ (152)
Since relation (152) determines $B^{\ast}$ uniquely, this concludes the proof. ∎
We conclude with
Proof of Theorem 2.24.
Let $\mathcal{H}$ and $\mathcal{K}$ be $p$-adic Hilbert spaces over the same quadratic extension $\mathbb{Q}_{p,\mu}$ of $\mathbb{Q}_{p}$, and let $\Phi\equiv\{\phi_{n}\}_{n\in\mathbb{N}}$ and $\Psi\equiv\{\psi_{m}\}_{m\in\mathbb{N}}$ be two orthonormal bases in $\mathcal{H}$ and $\mathcal{K}$ respectively. To prove that $\big\langle\mathcal{T}(\mathcal{H},\mathcal{K}),\|\cdot\|,\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}},\{\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\}_{j,k\in\mathbb{N}}\big\rangle$ is a $p$-adic Hilbert space, we first show that $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}$ is a (non-degenerate) $p$-adic inner product. Indeed, let $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$; we have:
$\langle\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}},T\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\mathrm{tr}\big(|\phi_{k}\rangle\langle\psi_{j}|T\big)=T_{jk},$ (153)
where, in the first equality, we used the fact that $\big(\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\big)^{\ast}=\langle\psi_{j},\cdot\hskip 1.13809pt\rangle\phi_{k}=|\phi_{k}\rangle\langle\psi_{j}|$. This then entails that
$\langle\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}},T\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=0,\;\;\forall j,k\in\mathbb{N}\quad\implies\quad T=0,$ (154)
i.e., $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}$ is non-degenerate. Moreover, by observing that
$\left|\langle S,T\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}\right|=\left|\mathrm{tr}(S^{\ast}T)\right|=\Big|\sum_{n}\langle\phi_{n},S^{\ast}T\phi_{n}\rangle\Big|\leq\|S\|\,\|T\|,$ (155)
we see that $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}$ satisfies the Cauchy-Schwarz inequality too, i.e., it is a non-Archimedean (or $p$-adic) inner product on $\mathcal{T}(\mathcal{H},\mathcal{K})$ (see Subsection 2.1). We now prove that $\{\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\}_{j,k\in\mathbb{N}}$ is an orthonormal basis in $\mathcal{T}(\mathcal{H},\mathcal{K})$. Indeed, it is clear that this set is orthonormal:
$\langle\hskip-1.99168pt\sideset{{}^{mn}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}},\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}=\mathrm{tr}\big(|\phi_{n}\rangle\langle\psi_{m}|\,|\psi_{j}\rangle\langle\phi_{k}|\big)=\delta_{mj}\delta_{nk},$ (156)
and, for every finite subset $I\subset\mathbb{N}\times\mathbb{N}$ and $\{\alpha_{ij}\}_{i,j\in I}\subset\mathbb{Q}_{p,\mu}$,
$\Big\|\sum_{i,j\in I}\alpha_{ij}\hskip-1.99168pt\sideset{{}^{ij}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\Big\|=\max_{i,j\in I}\left|\alpha_{ij}\right|.$ (157)
Let us now introduce the family of matrix operators $\sideset{{}^{l}}{}{\operatorname{\mathnormal{T}}}$, $l\in\mathbb{N}$, defined, for every trace class operator $T=\mathrm{op}_{\Phi,\Psi}(T_{mn})$, as
$\sideset{{}^{l}}{}{\operatorname{\mathnormal{T}}}\coloneqq\sum_{\max\{m,n\}\leq l}T_{mn}\hskip-1.99168pt\sideset{{}^{mn}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}.$ (158)
Then, $\forall l\in\mathbb{N}$, $\sideset{{}^{l}}{}{\operatorname{\mathnormal{T}}}\in\mathscr{B}_{\Phi,\Psi}$, and $\lim_{l}\sideset{{}^{l}}{}{\operatorname{\mathnormal{T}}}$ will converge to an operator in $\overline{\mathscr{B}_{\Phi,\Psi}}^{\|\cdot\|}(=\mathcal{T}(\mathcal{H},\mathcal{K}))$. We now want to prove that $\sideset{{}^{l}}{}{\operatorname{\mathnormal{T}}}$ converges precisely to $T$. Indeed, noting that
$\|T-\sideset{{}^{l}}{}{\operatorname{\mathnormal{T}}}\|=\Bigg\|\sum_{\max\{m,n\}>l}T_{mn}\hskip-1.99168pt\sideset{{}^{mn}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\Bigg\|=\sup\big\{\left|T_{mn}\right|\mid\max\{m,n\}>l\big\},$ (159)
and since $T\in\mathcal{T}(\mathcal{H},\mathcal{K})$ — which entails that, $\forall\epsilon>0$, there exists $j_{\epsilon}\in\mathbb{N}$ such that $\max\{m,n\}>j_{\epsilon}$ $\implies$ $\left|T_{mn}\right|<\epsilon$ — we have, for $l\geq j_{\epsilon}$
$\|T-\sideset{{}^{l}}{}{\operatorname{\mathnormal{T}}}\|=\sup\big\{\left|T_{mn}\right|\mid\max\{m,n\}>l\big\}<\sup\big\{\left|T_{mn}\right|\mid\max\{m,n\}>j_{\epsilon}\big\}<\epsilon,$ (160)
i.e., $\lim_{l}\sideset{{}^{l}}{}{\operatorname{\mathnormal{T}}}=T$. Concluding, $\{\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\}_{j,k\in\mathbb{N}}$ is an orthonormal system whose span is dense in $\mathcal{T}(\mathcal{H},\mathcal{K})$, i.e., it is an orthonormal basis in $\mathcal{T}(\mathcal{H},\mathcal{K})$. This concludes the proof as we have proven that the $p$-adic Banach space $\mathcal{T}(\mathcal{H},\mathcal{K})$ can be endowed with a $p$-adic inner product — i.e., $\langle\hskip 0.9105pt\cdot\hskip 1.70717pt,\cdot\hskip 0.9105pt\rangle_{\mbox{\tiny$\mathcal{T}(\mathcal{H},\mathcal{K})$}}$ — and the orthonormal basis $\{\hskip-1.99168pt\sideset{{}^{jk}\hskip-1.42262pt}{{}^{\Phi,\Psi}}{\operatorname{\mathnormal{E}}}\}_{j,k\in\mathbb{N}}$. ∎